Jefferson Lab HALL A Logbook for February 09 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

  • 263940 02/28/09 23:54 R. Michaels re: Qasy jumps.
  • 263937 02/28/09 23:37 Kalyan Cerenkov ADC spectrum shift
  • 263936 02/28/09 23:24 Kalyan preshower_E plot
  • 263935 02/28/09 23:06 Kalyan BigBite preshower_E vs shower_E
  • 263932 02/28/09 22:10 ycwang run 1906 and 20459 stopped for NMR Frequency Sweep
  • 263931 02/28/09 22:03 A. Rakhman Target screen snapshot
  • 263930 02/28/09 22:02 A. Rakhman Laser Absorption Spectrum
  • 263929 02/28/09 22:01 A. Rakhman NMR Frequency sweep (63.37 %)
  • 263924 02/28/09 21:55 Kalyan L1A check with BCM
  • 263921 02/28/09 21:01 ycwang plots not match with gold run
  • 263920 02/28/09 21:00 A. Rakhman Target Screen snapshot
  • 263917 02/28/09 20:31 A. Rakhman Target Position
  • 263916 02/28/09 20:25 bogdanw spot++
  • 263903 02/28/09 20:05 A. Rakhman Target Position
  • 263902 02/28/09 18:58 ycwang run 1904 and 20455 stopped because of no beam for half an hour
  • 263899 02/28/09 18:26 posik L1A for bb 1904
  • 263898 02/28/09 18:20 ycwang plots not match with gold run
  • 263897 02/28/09 18:02 A. Rakhman Target screen snapshot
  • 263896 02/28/09 18:00 A. Rakhman Laser Absorption Spectrum
  • 263895 02/28/09 17:59 A. Rakhman NMR Frequency sweep (63.39 %)
  • 263890 02/28/09 16:52 Kalyan BCM calibration : OLO2 vs Left HRS scaler
  • 263885 02/28/09 16:03 D. Parno / G Rate-Dependent Gain Shifts in Compton Photon Detector
  • 263883 02/28/09 16:01 Holmstrom Day Shift Summary
  • 263882 02/28/09 15:37 brads +1 MeV jump in beam energy for run 1900
  • 263879 02/28/09 15:19 Holmstrom Hall A Tools
  • 263876 02/28/09 14:24 Holmstrom Raster++ for 20453
  • 263873 02/28/09 14:10 Holmstrom Raster++
  • 263872 02/28/09 14:08 A. Deur NMR (64.4)
  • 263867 02/28/09 13:50 holmstrom harp scans
  • 263866 02/28/09 13:25 D. Parno L1A vs Clock Count for Run 1898
  • 263865 02/28/09 13:17 Holmstrom HARP Scan Problems
  • 263864 02/28/09 12:23 V. Sulkosky Experiment Summary for Feb 27, Owl - Feb 27, Swing
  • 263855 02/28/09 10:39 Holmstrom Half Wave plate switch
  • 263854 02/28/09 10:09 flay Compton CODA Restarted
  • 263853 02/28/09 09:59 flay Shift of Plots for run 20451
  • 263852 02/28/09 09:51 flay Gaps in Snyc Roc Plots (run 1898)
  • 263851 02/28/09 09:49 gjin routine MWDC check
  • 263847 02/28/09 08:42 Holmstrom MCC called RF problem fixed
  • 263846 02/28/09 08:26 Holmstrom MCC called
  • 263840 02/28/09 08:06 A. Deur Absorption Spectrum record
  • 263839 02/28/09 08:01 Fatiha Shift Summary
  • 263837 02/28/09 07:36 Fatiha Beam quality for the last 2h
  • 263834 02/28/09 07:21 Fatiha Started production with 3He, at KIn = I, P0 = 1.34,
  • 263833 02/28/09 07:21 hyao target status
  • 263832 02/28/09 07:17 hyao target move to he3
  • 263829 02/28/09 06:53 Fatiha Correction to the total charge
  • 263828 02/28/09 06:53 Fatiha Correction to the total charge
  • 263825 02/28/09 06:40 Fatiha Magnet screen at P0 = 1.34 GeV, Kin = I
  • 263824 02/28/09 06:36 hyao target moved to N2
  • 263823 02/28/09 06:36 Fatiha Moving to P0 = 1.340 GeV, kin = I
  • 263818 02/28/09 06:06 Fatiha Happex end run failure
  • 263815 02/28/09 05:34 Fatiha Total accumulated charge
  • 263814 02/28/09 05:03 Fatiha Better Beam quality
  • 263812 02/28/09 05:01 Fatiha Bigbite HV trips
  • 263800 02/28/09 04:06 Fatiha Spot++
  • 263797 02/28/09 03:58 Fatiha Beam is back
  • 263796 02/28/09 03:41 Fatiha Bad Bad beam quality
  • 263795 02/28/09 03:40 Fatiha MCC took beam to improve the quality
  • 263792 02/28/09 02:56 hyao target status
  • 263785 02/28/09 02:24 Fatiha Bad Beam quality
  • 263784 02/28/09 01:49 D. Parno Odd L1A for Run 1887
  • 263781 02/28/09 01:38 Fatiha Lecroy HV alarm
  • 263780 02/28/09 01:02 D. Parno Odd L1A for Run 1885
  • 263775 02/28/09 00:00 Ole Hansen Swing Shift Summary
  • Friday

  • 263769 02/27/09 23:24 Yunxiu Ye Abs_spectrum_022709_swing
  • 263768 02/27/09 23:15 Kalyan BCM calibration : 0L02 vs BCM1 and BCM2
  • 263767 02/27/09 23:02 D. Parno L1A vs clock count for Run 1855
  • 263766 02/27/09 22:51 D. Parno L1A vs clock count for Run 1858
  • 263765 02/27/09 22:40 D. Parno L1A vs clock count, Run 1859
  • 263764 02/27/09 22:38 D. Parno L1A vs clock count, Run 1861
  • 263763 02/27/09 22:36 D. Parno L1A vs Clock Count for Run 1875
  • 263760 02/27/09 22:27 Ole Hansen spot++
  • 263759 02/27/09 22:16 D. Parno Compton/HAPPEX restart
  • 263758 02/27/09 21:48 D. Parno Syncroc glitch in Run 1881
  • 263755 02/27/09 21:37 D. Parno bbps1:2002 S10, L10.0 turned back off after Run 1882
  • 263754 02/27/09 21:36 Yunxiu Ye Sweeping
  • 263740 02/27/09 19:52 Ole Hansen T6 threshold scan
  • 263735 02/27/09 19:19 A. Camsonne Compton electron detector run list
  • 263734 02/27/09 19:09 posik shift in histos for 1877
  • 263732 02/27/09 19:05 Ole Hansen Poor beam
  • 263731 02/27/09 19:04 Ole Hansen Compton HV trip
  • 263728 02/27/09 18:52 Re: T8 charge asymmetry
  • 263725 02/27/09 18:21 posik sync gaps in run 1875
  • 263715 02/27/09 17:12 Moffit Re: T8
  • 263711 02/27/09 16:39 Yunxiu Ye polarization measurement by NMR
  • 263709 02/27/09 16:30 Moffit Irregular helicity correlated T8.
  • 263707 02/27/09 16:02 Saha Shift Summary
  • 263704 02/27/09 15:56 flay Target Type for LHRS
  • 263697 02/27/09 14:06 flay Compton Reset
  • 263694 02/27/09 13:39 D. Parno Beam polarization estimate from Compton
  • 263693 02/27/09 12:37 Bo Zhao Screenshot for strip charts of target
  • 263692 02/27/09 12:35 Bo Zhao Screenshot for strip charts of target
  • 263691 02/27/09 12:33 Bo Zhao Laser Absorption Spectrum
  • 263690 02/27/09 12:30 Bo Zhao NMR Measurement
  • 263685 02/27/09 12:09 gjin routine MWDC check
  • 263668 02/27/09 09:31 V. Sulkosky Experiment Summary for Feb 26, Owl - Feb 26, Swing
  • 263665 02/27/09 08:34 Bo Zhao NMR Measurement
  • 263659 02/27/09 08:04 Fatiha Shift Summary
  • 263658 02/27/09 07:57 Fatiha Restarted auxiliary DAQs
  • 263655 02/27/09 07:34 dalton Sync Check1 run #1861
  • 263654 02/27/09 07:17 Fatiha HV for channel bbps1:2002 S10, L10.0 was off
  • 263653 02/27/09 07:13 Xinhu Yan Absorption spectrum (Owl)
  • 263648 02/27/09 06:59 adev BigBite WireChamber Measured Current change!
  • 263647 02/27/09 06:49 adev BigBite WireChamber Measured Current change!
  • 263646 02/27/09 06:34 dalton Restart BB coda
  • 263645 02/27/09 06:33 dalton Sync Check1 run #1859
  • 263640 02/27/09 05:30 Fatiha Beam current screen
  • 263639 02/27/09 05:20 dalton Sync Check run #1858
  • 263638 02/27/09 05:15 Faiha Prestarted auxiliary DAQs
  • 263637 02/27/09 05:02 Xinhu Yan NMR Frequency Sweep Screen Shot
  • 263622 02/27/09 04:03 dalton Sync Check run #1845
  • 263619 02/27/09 03:06 Fatiha Prestarted auxiliary DAQs
  • 263618 02/27/09 02:59 Fatiha Spot++
  • 263609 02/27/09 01:03 riordan Restarted Auxiliary DAQs
  • 263608 02/27/09 00:36 Xinhu Yan NMR Frequency Sweep Screen Shot
  • 263602 02/27/09 00:16 Ole Hansen Swing Shift Summary
  • Thursday

  • 263599 02/26/09 23:02 Ole Hansen halogscript diagnostics
  • 263598 02/26/09 22:49 Ole Hansen Disappearing halog directory
  • 263596 02/26/09 22:47 Ole Hansen Beam returns
  • 263594 02/26/09 22:45 Ole Hansen NMR
  • 263593 02/26/09 22:39 D. Parno Mystery HV alarm = Compton
  • 263592 02/26/09 22:31 riordan Spin state explanation
  • 263591 02/26/09 22:31 Jin Huang Re: Target Spin State Occasionall Reading 0
  • 263590 02/26/09 22:27 Ole Hansen Still no beam
  • 263589 02/26/09 22:27 Ole Hansen HV alarm fixed
  • 263588 02/26/09 22:26 Yunxiu Ye Abs_spectrum_022609_swing
  • 263586 02/26/09 22:22 Ole Hansen Mystery HV alarm
  • 263584 02/26/09 21:53 riordan Set Golden Runs
  • 263583 02/26/09 21:50 riordan Target Spin State Occasionall Reading 0
  • 263580 02/26/09 20:39 Yunxiu Ye NMR measurement
  • 263575 02/26/09 20:35 Yunxiu Ye Target miving
  • 263574 02/26/09 20:34 Ole Hansen 3He target
  • 263572 02/26/09 20:29 Ole Hansen N2 ref run kine K finished
  • 263560 02/26/09 20:00 Jack (Techs) Left Q3
  • 263559 02/26/09 19:56 Jack (Techs) Right Q3
  • 263555 02/26/09 19:35 Ole Hansen LHRS to 1.51 GeV/c
  • 263553 02/26/09 19:33 Ole Hansen BB replay woes (fixed)
  • 263552 02/26/09 19:06 Ole Hansen N2 reference runs kine M
  • 263546 02/26/09 18:50 Ole Hansen MWDC HV on
  • 263545 02/26/09 18:48 Ole Hansen spot++
  • 263544 02/26/09 18:45 V. Sulkosky Experiment Summary for Feb 25, Owl - Feb 25, Swing
  • 263543 02/26/09 18:45 Ole Hansen Rastersize check
  • 263541 02/26/09 18:30 Yunxiu Ye target moving
  • 263540 02/26/09 18:27 Yunxiu Ye Target moving
  • 263539 02/26/09 18:26 Ole Hansen Access done. Prep for ref cell runs
  • 263538 02/26/09 18:24 brads BB Cer PMT 18 TDC cable wasn't seated well -- fixed
  • 263536 02/26/09 18:01 Ole Hansen Controlled access
  • 263535 02/26/09 17:59 Ole Hansen spot++
  • 263534 02/26/09 17:56 D. Parno BB Cerenkov TDC 18 is missing
  • 263530 02/26/09 17:44 brads BB Cerenkov gains checked -- RUNLIST
  • 263527 02/26/09 17:32 Ole Hansen Harp scans
  • 263523 02/26/09 17:20 D. Parno Syncroc check -- Run 1810
  • 263519 02/26/09 17:00 Yunxiu Ye Target moving
  • 263517 02/26/09 16:37 Kalyan T6 threshold changed to -110mV
  • 263514 02/26/09 16:10 Higinbotham Left Arm Quads Cycled
  • 263512 02/26/09 16:07 V. Sulkosky Turned off MWDC HV
  • 263511 02/26/09 16:05 Higinbotham Shift Summary
  • 263509 02/26/09 16:04 A. Camsonne Compton Edet Run plan D test strip 0 to 95 16074
  • 263508 02/26/09 15:54 F. Garibaold absorption sepctra day-26-February
  • 263507 02/26/09 15:51 A. Camsonne Compton Edet Run plan C test strip 0 to 95 16073
  • 263506 02/26/09 15:41 A. Camsonne Compton Edet Run plan C test strip 96 to 191 16072
  • 263505 02/26/09 15:33 A. Camsonne Compton Edet Run plan B test strip 96 to 191 16071
  • 263504 02/26/09 15:18 A. Camsonne Compton Edet Run plan B test strip 0 to 95 16070
  • 263503 02/26/09 15:12 A. Camsonne Test Compton Edet plan A
  • 263502 02/26/09 14:41 A. Camsonne Compton Edet run 16069
  • 263501 02/26/09 14:17 A. Camsonne run fom strip 0 to 95 Plan A
  • 263500 02/26/09 13:38 A. Camsonne Run 16068 compton edet test
  • 263499 02/26/09 13:25 brads Old ADC filter from preshower is in cabinet
  • 263487 02/26/09 12:29 Kalyan shower channels 19.4 , 22.7 checked - looks okay
  • 263472 02/26/09 11:53 A. Camsonne Power cycled network switched in right HRS
  • 263471 02/26/09 11:36 Jin Huang Target spin direction rotated to 270 degrees; All Back to Normal d2n Settings
  • 263466 02/26/09 11:23 Kalyan BigBite high voltages updated
  • 263460 02/26/09 10:48 ellie Target changed during Run 1816
  • 263456 02/26/09 10:42 gjin routine MWDC check
  • 263455 02/26/09 10:41 Kalyan Re: shower holes known channels
  • 263454 02/26/09 10:33 brads BB Cerenkov HV off on PMT 11--20
  • 263451 02/26/09 09:02 brads BB Threshold set to -20mV for cosmics
  • 263450 02/26/09 08:36 Garibaldi Target auto spin flip off
  • 263449 02/26/09 08:33 cusanno range of reference plots for LHS
  • 263448 02/26/09 08:29 cusanno Pion Rejector E/P run 20422
  • 263439 02/26/09 07:41 dalton screenshot of Sync Check1 run #1822
  • 263434 02/26/09 07:03 dalton BB Cer ADC offset again
  • 263431 02/26/09 06:55 Fatiha BCM screen for the last 2h ( beam down twice...)
  • 263430 02/26/09 06:19 dalton screenshot of BB Cer ADCs for Run #1819
  • 263429 02/26/09 06:16 Fatiha Beam is back
  • 263424 02/26/09 06:00 Fatiha MCC took beam for some optics test
  • 263415 02/26/09 04:53 dalton Sync Check1 run 1818
  • 263400 02/26/09 04:04 dalton BB prescale changed
  • 263387 02/26/09 02:00 Fatiha Megnet Screen at P0=0.800 GeV
  • 263385 02/26/09 01:48 Fatiha Setting P0 to 0.800 GeV
  • 263383 02/26/09 01:25 dalton BB_Sh to -0.110 V
  • 263382 02/26/09 01:24 xqian Holes in the shower.
  • 263381 02/26/09 01:23 xqian database update for 20090225 for the BigBite Shower Counter
  • 263376 02/26/09 01:15 Xinhu Yan Target changed to Reference cell
  • 263371 02/26/09 00:49 Fatiha Spot++
  • Wednesday

  • 263365 02/25/09 23:49 R. Subedi Shift Summary (Swing)
  • 263364 02/25/09 23:19 brads Live camera views available on hacweb1 (in Counting House only, please)
  • 263359 02/25/09 22:40 xqian RUN LIST for trigger alignment
  • 263357 02/25/09 22:37 R. Subedi Hall A Counting House TV monitor dies [Techs]
  • 263356 02/25/09 22:37 Kalyan Corrected HV file for bigbite
  • 263346 02/25/09 21:46 Jin Huang Correction Coil -> 0.9A
  • 263340 02/25/09 20:41 xqian RUN LOG for Monday pulser test for the BigBite Trigger system
  • 263337 02/25/09 20:35 Jin Huang Correction Coil -> 0.8A
  • 263336 02/25/09 20:22 R. Subedi Spot check
  • 263315 02/25/09 18:58 Kalyan BigBite HV change
  • 263314 02/25/09 18:56 D. Parno Syncroc check: Run 1797
  • 263313 02/25/09 18:55 mcanan sync check1 screen shot
  • 263312 02/25/09 18:54 V. Sulkosky Experiment Summary for Feb 23, Swing - Feb 24, Swing
  • 263310 02/25/09 18:27 D. Parno HAPPEX/Compton DAQs restarted
  • 263303 02/25/09 17:35 D. Parno L1A vs clock count for Run 1787
  • 263302 02/25/09 17:32 D. Parno L1A vs clock count for Run 1784
  • 263301 02/25/09 17:29 D. Parno L1A vs clock count for Run 1795
  • 263300 02/25/09 17:23 D. Parno LHRS good electron rate for 4K kinematic
  • 263299 02/25/09 17:19 mcanan Sync Check
  • 263298 02/25/09 16:59 D. Parno Good electron rate for 4J kinematic
  • 263294 02/25/09 16:38 mcanan BB Deat Time>20%
  • 263293 02/25/09 16:36 Kalyan ROC 5 errors
  • 263289 02/25/09 16:31 R. Subedi BigBite DAQ Deadtim 20%
  • 263288 02/25/09 16:20 A. Deur Laser Absorption Spectrum
  • 263287 02/25/09 16:05 D. Parno Good Compton spectrum after this morning's high background
  • 263286 02/25/09 16:05 Jin Huang Correction Coil -> 1.5A
  • 263285 02/25/09 16:02 Higinbotham Shift Summary
  • 263283 02/25/09 15:57 Higinbotham VESDA Levels Down Significantly
  • 263282 02/25/09 15:54 gjin routine MWDC check
  • 263279 02/25/09 15:16 Bo Zhao Target control system report
  • 263276 02/25/09 14:59 Higinbotham Left arm ROCs rebooted
  • 263275 02/25/09 14:57 X. Zheng error message in ROC5
  • 263272 02/25/09 14:48 Higinbotham Hall A Tools
  • 263270 02/25/09 14:42 Higinbotham raster++ (a.k.a. spot++)
  • 263263 02/25/09 12:43 A. Camsonne Compton detector photon rates
  • 263262 02/25/09 12:30 Higinbotham Hall A Current Montior
  • 263261 02/25/09 12:25 Bo Zhao Screenshot for strip charts of target
  • 263260 02/25/09 12:21 Bo Zhao Laser Absorption Spectrum
  • 263259 02/25/09 12:07 X. Zheng BB sync check run 1788
  • 263250 02/25/09 11:45 Higinbotham Access Completed (VESDA Picture Attached)
  • 263245 02/25/09 08:56 Higinbotham Disk Replaced (late entry)
  • 263244 02/25/09 08:43 Higinbotham VESDA Check @ 9:15
  • 263243 02/25/09 08:40 A. Deur gap in sync check for run 1787
  • 263240 02/25/09 08:03 Fatiha Shift Summary
  • 263236 02/25/09 07:53 dalton VESDA: Folts and Kaush informed
  • 263235 02/25/09 07:47 Fatiha VESDA Alerts
  • 263228 02/25/09 05:23 Fatiha Spot++
  • 263227 02/25/09 05:14 dalton Target Laser Spectrum
  • 263224 02/25/09 04:31 Muangma sync Check run bb 1784
  • 263209 02/25/09 02:18 D. Parno L1A vs Clock for #1782
  • 263208 02/25/09 02:03 Muangma sync check run BB1782
  • 263207 02/25/09 01:50 D. Parno LHRS VDC Efficiency
  • 263198 02/25/09 00:22 D. Parno LHRS golden run set to 20392
  • 263196 02/25/09 00:16 jpchen Swing shift summary
  • 263195 02/25/09 00:04 Jin Huang Correction Coil On to Reduce Massing
  • Tuesday

  • 263192 02/24/09 22:54 jpchen spot++
  • 263189 02/24/09 22:48 dwang target moved back to he-3
  • 263184 02/24/09 22:12 deng target moved to reference cell
  • 263180 02/24/09 22:08 jpchen LHRS p0 to 1.51 GeV/tool screen
  • 263160 02/24/09 18:42 jpchen Compton HV tripped. Reset.
  • 263155 02/24/09 18:06 jpchen Changed the BCM calibration constants to the correct values
  • 263154 02/24/09 17:55 jpchen icoha5 (for BCMs) rebooted
  • 263153 02/24/09 16:58 A. Camsonne Compton Edet test run 16024 16025
  • 263150 02/24/09 16:39 Franklin/Par Compton FADC problem "fixed"
  • 263149 02/24/09 16:11 brads Mystery HV alarm
  • 263148 02/24/09 16:07 Higinbotham Shift Summary
  • 263144 02/24/09 15:49 gjin routine MWDC check
  • 263143 02/24/09 15:44 A. Camsonne Background rates
  • 263140 02/24/09 15:36 A. Camsonne Compton GSO PMT tripped
  • 263139 02/24/09 15:28 Higinbotham LeCroy HV Alarm
  • 263137 02/24/09 14:54 Higinbotham Coda back to bigbite_buf mode
  • 263135 02/24/09 14:47 A. Camsonne Background rates in Compton
  • 263134 02/24/09 14:37 fansler VESDA Alarms
  • 263133 02/24/09 14:35 yawei L1A vs clock count for run#1771
  • 263132 02/24/09 14:33 yawei L1A vs clock count for run#1771
  • 263131 02/24/09 14:30 Higinbotham BigBite Prescale
  • 263130 02/24/09 14:01 Saha Gaps in Sync Check1 in BB1771
  • 263127 02/24/09 13:52 ellie L1A vs Clock Count for Run 1770
  • 263124 02/24/09 13:29 Ole Hansen adaql2 bus speeds
  • 263123 02/24/09 13:13 Higinbotham Hall A Tools
  • 263122 02/24/09 13:11 Saha Gaps in Sync Check in BB1770
  • 263121 02/24/09 13:10 Higinbotham Raster++
  • 263116 02/24/09 12:12 Kalyan missing shower channel 28 ADC fixed
  • 263115 02/24/09 12:10 Higinbotham Shower Cable Fixed
  • 263106 02/24/09 11:54 Jack VDC gas shut off
  • 263105 02/24/09 11:34 ellie Gaps in Sync Check for Run 1768
  • 263104 02/24/09 11:31 Saha Gaps in Sync Check1 (st. lines) in BB1768
  • 263103 02/24/09 11:28 Higinbotham North Linac Problem
  • 263102 02/24/09 11:27 shahinyan target
  • 263101 02/24/09 11:25 A. Camsonne Compton Edet test run 16015 16016
  • 263100 02/24/09 11:22 ellie Left Arm VDC Efficiency Capture
  • 263095 02/24/09 10:37 ellie L1A vs Clock Count for Run 1757
  • 263094 02/24/09 10:06 ellie L1A vs Clock Count for Run 1760
  • 263093 02/24/09 09:59 ellie L1A vs Clock Count for Run 1761
  • 263092 02/24/09 09:59 folts tech/ elevated VESDA
  • 263089 02/24/09 09:44 Higinbotham Beam Back
  • 263088 02/24/09 09:44 Jack RICH status
  • 263087 02/24/09 09:32 ellie L1A vs Clock Count for Run 1766
  • 263086 02/24/09 09:27 Higinbotham Hall A Vesda Alert (Controlled Access)
  • 263085 02/24/09 09:20 Saha Gaps in Sync Check (st. lines) in BB1766
  • 263082 02/24/09 09:14 R. Michaels re: disks full
  • 263081 02/24/09 09:03 Higinbotham Coda Changed To Flash ADC Mode
  • 263080 02/24/09 09:00 V. Sulkosky Deadtime Hiccups for run 1766
  • 263079 02/24/09 08:58 ellie Re: Compton DAQ rates high
  • 263076 02/24/09 08:44 ellie Compton DAQ rates high
  • 263073 02/24/09 08:02 Fatiha Shift Summary
  • 263065 02/24/09 06:49 Fatiha Hall A Beam Charge Asymmetry
  • 263064 02/24/09 06:20 cusanno data3/ and data4/ disks rather full
  • 263059 02/24/09 06:01 cusanno Dead Time on BB
  • 263058 02/24/09 05:59 cusanno ROC sync failure, run 1761
  • 263057 02/24/09 05:56 Fatiha Compton RATE_G not working?
  • 263054 02/24/09 05:14 cusanno ROC sync failure, run 1760
  • 263049 02/24/09 04:19 Fatiha Spot++
  • 263048 02/24/09 04:11 dalton Polarized He3 Target Control System Status Report
  • 263047 02/24/09 04:09 dalton Target Laser Spectrum
  • 263032 02/24/09 02:57 cusanno roc sync lost on run 1757
  • 263023 02/24/09 00:38 dalton VESDA Hall A alert
  • 263022 02/24/09 00:03 deng inconsistence of Sync check for BB run 1753
  • 263021 02/24/09 00:03 F. Garibaldi swing shift
  • Monday

  • 263013 02/23/09 22:58 V. Sulkosky Re: strange s1 tdc?
  • 263010 02/23/09 22:33 Jin Huang BB Golden Run -> 1748
  • 263009 02/23/09 22:28 deng deadtime more than 20% in BB run 1751
  • 263006 02/23/09 22:12 V. Sulkosky Experiment Summary for Feb 22, Swing - Feb 23, Day
  • 263005 02/23/09 22:12 xqian Shower channel 2.2 is missing
  • 263003 02/23/09 22:04 deng wire chamber trips for BB run 1750
  • 263001 02/23/09 21:54 xqian strange s1 tdc?
  • 263000 02/23/09 21:51 Moffit HAPPEX DAQ: d2n Config
  • 262999 02/23/09 21:38 deng inconsistence for sync check
  • 262998 02/23/09 21:29 xqian We will do the change tomorrow 02-24
  • 262995 02/23/09 21:04 R. Michaels evtype 120 reappearing (L-arm)
  • 262990 02/23/09 20:11 D. Parno Offline deadtime < 20% for PS2 = 1
  • 262989 02/23/09 20:05 D. Parno LHRS golden run set to 20373
  • 262980 02/23/09 19:52 xqian New database for the shower coutner adjustment
  • 262979 02/23/09 19:52 A. Camsonne Compton Edet test run 16002
  • 262978 02/23/09 19:51 A. Camsonne Compton Edet test run 16001
  • 262977 02/23/09 19:45 D. Parno Synchronization checks (L1A vs. clock count)
  • 262972 02/23/09 19:18 A. Camsonne Compton Edet test run 15998
  • 262971 02/23/09 19:18 A. Camsonne Compton Edet test run 15997
  • 262970 02/23/09 19:13 dwang target moved to He3
  • 262967 02/23/09 18:59 A. Camsonne Compton Edet test run 15996
  • 262966 02/23/09 18:49 V. Sulkosky Re-downloaded Coda for Left HRS
  • 262961 02/23/09 18:35 V. Sulkosky BigBite MWDC HV On
  • 262960 02/23/09 18:34 deng rate check for different raster sizes
  • 262957 02/23/09 18:31 D. Parno Syncroc studies: L1A vs Clock for Run 1649
  • 262956 02/23/09 18:28 R. Michaels re: L-HRS runstart problem (fixed)
  • 262954 02/23/09 18:25 deng spot++ check 6*6
  • 262951 02/23/09 18:18 deng spot++ check 4*4
  • 262945 02/23/09 17:53 Jin Huang Spin Signal on LHRS is OK
  • 262942 02/23/09 17:43 D. Parno Tuned Compton cavity
  • 262940 02/23/09 17:41 V. Sulkosky Raster Check (2 x 2.6) on Reference Cell
  • 262935 02/23/09 17:32 V. Sulkosky Spot++ 2 x 2.6 (MCC units)
  • 262934 02/23/09 17:30 Kalyan high voltage adjusted for preshower left 1-16 channel
  • 262933 02/23/09 17:27 deng spot++ check
  • 262932 02/23/09 17:27 Jin Huang Spin Flip Started for 4pass Single Spin Asym running
  • 262930 02/23/09 17:21 D. Parno T6 threshold set to -133 mV
  • 262928 02/23/09 17:05 dwang target moved to optics
  • 262927 02/23/09 17:05 Kalyan Low pass filter changed on preshower left ADC (1-16) channels
  • 262925 02/23/09 16:00 jpchen Day shift summary
  • 262924 02/23/09 15:58 jpchen 114 psig N2 in reference cell
  • 262923 02/23/09 15:57 V. Sulkosky MWDC HV Off for Tuning
  • 262921 02/23/09 15:52 jpchen Target to Empty
  • 262912 02/23/09 15:01 gjin MWDC check (graphs)
  • 262911 02/23/09 15:00 ge jin routine MWDC check
  • 262910 02/23/09 14:57 D. Parno Syncroc studies: L1A vs Clock for Run 1688
  • 262909 02/23/09 14:55 jpchen MCC called: 1/2 hour heads up for hall be in beam permit
  • 262906 02/23/09 14:52 D. Parno Syncroc studies: L1A vs Clock for Run 1662
  • 262905 02/23/09 14:40 D. Parno Syncroc studies: L1A vs Clock for Run 1647
  • 262901 02/23/09 14:31 D. Parno Syncroc studies: L1A vs Clock for Run 1603
  • 262900 02/23/09 14:19 D. Parno Syncroc studies: L1A vs Clock for Run 1597
  • 262896 02/23/09 13:55 A. Camsonne Compton test run 15976
  • 262895 02/23/09 13:32 A. Camsonne Compton test run 15975
  • 262894 02/23/09 13:31 A. Camsonne Compton test run 15974
  • 262893 02/23/09 13:29 A. Camsonne Compton test run 15973
  • 262892 02/23/09 13:28 A. Camsonne Compton test run 15972
  • 262891 02/23/09 13:27 A. Camsonne Compton test run 15971
  • 262890 02/23/09 13:24 Parno / Fran New PMT/Base for Compton Photon Detector
  • 262886 02/23/09 13:13 Jin/Yi Field Rotation to Trans-; Broken fiber; Power test
  • 262883 02/23/09 13:06 Yi/Jin Downstream Target Chamber Pickup coil solved
  • 262876 02/23/09 12:25 jpchen Tools screen
  • 262875 02/23/09 12:19 jpchen LHRS P changed from 0.9 GeV to 1.6 GeV
  • 262874 02/23/09 12:16 jpchen reference system gauges re-calibrated.
  • 262865 02/23/09 11:50 Jin/Yi Internal Temperature Test for Long. Pumping
  • 262859 02/23/09 11:39 brads re: BB Scint TDCs shifted (fixed via ROC5 power cycle)
  • 262858 02/23/09 11:38 Parno / Fank Compton beam pipe heater unplugged
  • 262850 02/23/09 11:26 brads re: BB Scint TDCs shifted (Run 1719)
  • 262846 02/23/09 10:53 Yi/Jin NMR Cross Calibration for Vertical pumping
  • 262845 02/23/09 10:47 Yi/Jin EPR Calibration with Longitudinal Pumping
  • 262842 02/23/09 10:09 ellie BB Scint peaks still wide for Cosmics (Run 1722)
  • 262841 02/23/09 10:08 ellie TS12 Rebooted
  • 262836 02/23/09 09:52 R. Michaels Target NMR and movements.
  • 262831 02/23/09 09:31 ellie BB Shower T6 Threshold set to -20mV
  • 262830 02/23/09 09:12 ellie Slight shift in Sync Check2 for Run 1719
  • 262829 02/23/09 09:08 ellie Run 1719 BB Scint - TSum time
  • 262828 02/23/09 09:07 jpchen Beam off for pass change/RF recovery, controleld access
  • 262827 02/23/09 09:06 ellie BB Scint TDCs shifted (Run 1719)
  • 262824 02/23/09 08:50 golge BB Run #1719 shift in BB Cher ADC
  • 262823 02/23/09 08:49 ellie CODA crashed during 1718
  • 262822 02/23/09 08:47 R. Michaels data disk cleanup (e06010)
  • 262821 02/23/09 08:33 ellie Sync Check (Runs 1711, 1712)
  • 262815 02/23/09 08:01 Ibrahim Owl Shift Summary (Feb. 23, 2009)
  • 262807 02/23/09 06:52 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262804 02/23/09 06:29 Ibrahim Spot++ for Run 20355
  • 262803 02/23/09 06:28 dalton Target moved to He3 Cell
  • 262801 02/23/09 06:25 dalton Target moved to Optics
  • 262800 02/23/09 06:23 Ibrahim Beam Optics Check
  • 262798 02/23/09 06:13 dalton Target Laser Spectrum
  • 262797 02/23/09 06:13 dalton Target stripcharts
  • 262796 02/23/09 06:12 dalton Target Laser Spectrum
  • 262795 02/23/09 06:07 dalton Target moved to empty position
  • 262794 02/23/09 06:06 dalton Target: NMR Frequency sweep (56.51 %)
  • 262785 02/23/09 04:26 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262774 02/23/09 03:54 Ibrahim Spot++ for Run 20352
  • 262765 02/23/09 02:06 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262764 02/23/09 01:59 Ibrahim Half-wave Plate is OUT
  • 262763 02/23/09 01:59 dalton Target stripcharts
  • 262762 02/23/09 01:57 dalton Target Laser Spectrum
  • 262759 02/23/09 01:55 dalton Target: NMR Frequency sweep (56.86 %)
  • 262758 02/23/09 01:49 brads Expert/3rd crew: please watch for and log any sync check gaps
  • 262757 02/23/09 01:49 dalton Target Laser Spectrum
  • 262747 02/23/09 00:02 flay Shift Summary (Swing) for 2/22/09
  • Sunday

  • 262742 02/22/09 23:07 deng screen shot for NMR
  • 262725 02/22/09 20:19 brads Compton cavity power dropping?
  • 262716 02/22/09 19:54 V. Sulkosky Experiment Summary for Feb 18, Swing - Feb 19, Swing
  • 262711 02/22/09 19:12 V. Sulkosky Experiment Summary for Feb 21, Day - Feb 22, Day
  • 262710 02/22/09 18:33 Yi Qiang Spin direction rotated to 0 degree, polarization: 61.86%
  • 262704 02/22/09 17:26 deng NMR frequency sweep
  • 262702 02/22/09 17:23 bogdanw regarding GC analysis
  • 262701 02/22/09 17:22 A. Camsonne Compton electron detector currents
  • 262700 02/22/09 17:21 R. Subedi Spot check on 3He now OK
  • 262699 02/22/09 17:20 brads Next run's raster looks OK
  • 262697 02/22/09 17:15 brads/subedi Strange raster pattern?
  • 262693 02/22/09 16:59 deng target moved to He3
  • 262691 02/22/09 16:57 deng raster check
  • 262685 02/22/09 16:37 zhangyi Day shift summary
  • 262681 02/22/09 15:17 A. Camsonne Compton run 15937
  • 262680 02/22/09 15:16 A. Camsonne Compton un 15936
  • 262679 02/22/09 15:14 A. Camsonne Compton Edet test run 15935
  • 262678 02/22/09 14:08 gjin routine MWDC check
  • 262677 02/22/09 13:59 Yi Qiang Gas Cherenkov study from ep elastics data run 1258-1262
  • 262675 02/22/09 13:23 posik BB gas cer ped run 1699
  • 262671 02/22/09 13:02 posik golden run->20332,1650
  • 262670 02/22/09 12:52 Zhihong Ye Target moved to Optcs(BeO)
  • 262669 02/22/09 12:51 Zhihong Ye 2nd_NMR_02_22_day_shift
  • 262668 02/22/09 12:50 Zhihong Ye Target screen shot -2
  • 262651 02/22/09 11:30 Zhihong Ye Target moved to He3
  • 262646 02/22/09 10:52 Zhihong Ye Target moved to refference cell N2
  • 262645 02/22/09 10:48 Zhihong Ye Fill N2 with 100 psig
  • 262642 02/22/09 10:24 posik LHRS coda died run 20329 junk
  • 262636 02/22/09 08:51 Zhihong Ye Screen-shot for Target (2)
  • 262634 02/22/09 08:49 Zhihong Ye 1st_NMR_02_22_day_shift
  • 262628 02/22/09 08:03 Ibrahim Owl Shift Summary (Feb. 22, 2009)
  • 262624 02/22/09 07:32 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262613 02/22/09 05:35 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262612 02/22/09 05:10 hyao Sync check2 problem
  • 262611 02/22/09 04:54 dalton Target stripcharts
  • 262610 02/22/09 04:52 dalton Laser Absorption Spectrum
  • 262607 02/22/09 04:51 dalton NMR Frequency sweep (61.73 %)
  • 262604 02/22/09 04:10 Ibrahim Spot++ for Run 20324
  • 262591 02/22/09 03:05 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262586 02/22/09 02:33 hyao sync check 2 problem
  • 262585 02/22/09 02:14 hyao wrong title of Left HRS onl-analysics figure
  • 262582 02/22/09 02:07 gjin only 1 sync graph has problem during runs 1675-1684
  • 262577 02/22/09 00:36 dalton NMR frequency sweep (61.78 %)
  • 262576 02/22/09 00:18 dalton Absorption Spectrum
  • 262570 02/22/09 00:05 R. Subedi Shift Summary (Swing)
  • Saturday

  • 262569 02/21/09 23:27 ge jin problematic sync graphs
  • 262558 02/21/09 21:25 R. Subedi Spot check for checklist
  • 262557 02/21/09 20:27 A. Rakhman Target screen snapshots
  • 262556 02/21/09 20:27 R. Subedi Beam back
  • 262555 02/21/09 20:27 A. Rakhman Laser Absorption Spectrum
  • 262554 02/21/09 20:25 A. Rakhman NMR Frequency sweep (62.03 %)
  • 262551 02/21/09 20:19 R. Subedi Beam down due to ARC4 magnet problem
  • 262538 02/21/09 19:41 V. Sulkosky Re: Experiment Summary for Feb 17, Day - Feb 18, Day
  • 262537 02/21/09 19:39 posik rate for kinematic F90
  • 262534 02/21/09 19:27 V. Sulkosky Experiment Summary for Feb 17, Day - Feb 18, Day
  • 262533 02/21/09 18:59 posik vnc server terminate
  • 262530 02/21/09 18:28 brads re: rsync speed
  • 262525 02/21/09 17:42 V. Sulkosky Experiment Summary for Feb 20, Owl - Feb 21, Owl
  • 262524 02/21/09 17:34 R. Michaels re: rsync speed
  • 262523 02/21/09 17:20 R. Michaels replace rcp by rsync
  • 262522 02/21/09 17:06 A. Rakhman Laser Absorption Spectrum
  • 262521 02/21/09 17:05 A. Rakhman Target screen snapshots
  • 262520 02/21/09 17:03 A. Rakhman NMR Frequency sweep (61.94 %)
  • 262513 02/21/09 16:17 posik golden runs-> 20314,1650
  • 262512 02/21/09 16:01 Holmstrom Day Shift Summary
  • 262508 02/21/09 15:35 A. Deur Pol He3 cell in beam position
  • 262505 02/21/09 15:04 golge BB #1672 HRS #20313 is N2 not He3
  • 262502 02/21/09 14:50 A. Deur ref cell in beam position
  • 262501 02/21/09 14:50 A. Deur NMR Measurement (61.89%)
  • 262500 02/21/09 14:48 D. Parno HAPPEX/Compton restart
  • 262499 02/21/09 14:46 A. Deur ref cell filled with N2
  • 262497 02/21/09 14:40 posik BB sync gap
  • 262489 02/21/09 11:59 golge BB Run #1669 shift in BB Cher ADC plot
  • 262472 02/21/09 09:25 A. Deur Target polarisation is 61.52%
  • 262469 02/21/09 08:41 R. Michaels backing up Compton FADC data again
  • 262466 02/21/09 08:12 A. Deur Absorption Spectrum
  • 262465 02/21/09 08:09 A. Deur Record of target temperature, holding field, etc...
  • 262463 02/21/09 08:02 Ibrahim Owl Shift Summary (Feb. 21, 2009)
  • 262460 02/21/09 07:31 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262451 02/21/09 05:22 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262450 02/21/09 05:05 Ibrahim Sync Check2 Gaps
  • 262449 02/21/09 04:51 Nelyubin Holding Field
  • 262448 02/21/09 04:50 Nelyubin Oven temperature
  • 262447 02/21/09 04:47 Nelyubin Absorption Spectrum
  • 262446 02/21/09 04:42 Nelyubin NMR sweep owl shift 02/21/09 4:35 polarization 62.24% transv +
  • 262429 02/21/09 02:54 Ibrahim Spot++ for Run 20305
  • 262426 02/21/09 02:40 D. Parno HAPPEX/Compton runs restarted
  • 262425 02/21/09 02:38 Ibrahim MCC Magnet Trip
  • 262424 02/21/09 02:36 brads Restarted auto-snap software
  • 262421 02/21/09 01:51 D. Parno BB prescale, configuration changed; crates reset
  • 262410 02/21/09 00:12 Kalyan Re:BB Run 1654 High Deadtime
  • 262406 02/21/09 00:00 Ole Hansen Swing Shift Summary
  • Friday

  • 262405 02/20/09 23:59 Nelyubin NMR Sweep Shot , polarization 62% transverse +
  • 262404 02/20/09 23:57 flay BB Run 1654 High Deadtime
  • 262403 02/20/09 23:53 A. Camsonne Run Edet test compton 15917 Calibration
  • 262402 02/20/09 23:53 Ole Hansen Target NMR
  • 262399 02/20/09 23:49 A. Camsonne Run Edet test compton 15916 Calibration
  • 262398 02/20/09 23:42 D. Parno Compton/HAPPEX runs restarted
  • 262397 02/20/09 23:36 A. Camsonne Run Edet test compton starting at 15903
  • 262394 02/20/09 23:27 flay Golden ROOTfiles set to 20299, 1650
  • 262393 02/20/09 23:23 Ole Hansen bb_buf_FADC mode (late entry)
  • 262392 02/20/09 23:16 D. Parno Compton/Moller Comparison
  • 262391 02/20/09 23:14 Ole Hansen spot++
  • 262388 02/20/09 22:39 Nelyubin Absorption line spectrum
  • 262387 02/20/09 22:35 Nelyubin Strip Charts holding field and polarization
  • 262386 02/20/09 22:33 adaq STRIP CHARTS SHOT
  • 262383 02/20/09 21:26 Ole Hansen More /adaqfs/home cleanup
  • 262381 02/20/09 21:10 Ole Hansen BB HV trip
  • 262377 02/20/09 20:40 flay Run 1651 is a junk run
  • 262372 02/20/09 19:54 posik Fig. for H kinematic good electrons
  • 262365 02/20/09 19:31 posik sync gaps in run 1647
  • 262364 02/20/09 19:13 Yi Qiang Target Spin Rotated to 90 degrees: Transverse -, 60.36%
  • 262359 02/20/09 18:16 Ole Hansen Production resumes
  • 262356 02/20/09 18:05 flay Cannot run analyzer on adaql4,5,6,7
  • 262355 02/20/09 18:05 flay Cannot run analyzer on adaql4,5,6,7
  • 262352 02/20/09 17:54 Ole Hansen /adaqfs/home full - cleanup
  • 262351 02/20/09 17:42 flay Stack Trace error on run 1644/20294
  • 262350 02/20/09 17:38 Ole Hansen Beam trips
  • 262349 02/20/09 17:36 Ole Hansen HV trip
  • 262348 02/20/09 17:34 Ole Hansen HV alarms
  • 262347 02/20/09 17:14 flay Golden Root files set to 20291, 1637
  • 262346 02/20/09 17:06 Kalyan Blaster test - network speed looks good
  • 262344 02/20/09 16:47 brads re: ROC 5 , 792 ADC errors
  • 262342 02/20/09 16:41 Kalyan ROC 5 , 792 ADC errors
  • 262338 02/20/09 16:25 flay Golden Root file for LHRS set to 20289
  • 262337 02/20/09 16:17 Kalyan messages in TS1 terminal ...
  • 262335 02/20/09 16:02 X. Zheng Shift summary (day)
  • 262334 02/20/09 15:52 adaq routinely MWDC check
  • 262333 02/20/09 15:51 Kalyan Level 1 accept is fluctuating accept is fluctuating
  • 262328 02/20/09 15:25 X. Zheng spot++ (graph)
  • 262327 02/20/09 15:24 Nuruzzaman Energy Transfer shifted for Lt.HRS
  • 262326 02/20/09 15:24 X. Zheng spot++ run 20291
  • 262325 02/20/09 15:23 X. Zheng 15uA
  • 262312 02/20/09 14:58 Kalyan Preshower sum input line swapped for left and right
  • 262309 02/20/09 14:43 LeRose Changed Target to 3He
  • 262308 02/20/09 14:31 LeRose Target changed Reference Cell
  • 262307 02/20/09 14:28 A. Camsonne Compton run test 15899 and 15900
  • 262304 02/20/09 14:27 X. Zheng spot++ on optics target
  • 262303 02/20/09 14:26 X. Zheng spot++ on optics target
  • 262300 02/20/09 14:16 LeRose NMR Frequency sweep
  • 262299 02/20/09 14:14 LeRose Target changed Optics
  • 262295 02/20/09 14:05 brads Moved bbsfi2, bbsfi3 to their own network switch & port
  • 262289 02/20/09 13:38 flay Compton Rebooted
  • 262288 02/20/09 13:34 X. Zheng Controlled access
  • 262285 02/20/09 12:59 X. Zheng spot++ run 20289
  • 262284 02/20/09 12:56 LeRose Absorption Spectrum
  • 262271 02/20/09 10:57 A. Camsonne Run 15895 Edet run compton
  • 262270 02/20/09 10:51 A. Camsonne Run 15893 Edet run compton
  • 262269 02/20/09 10:32 A. Camsonne Edet compton run 15891
  • 262268 02/20/09 10:23 A. Camsonne Edet compton run 15889
  • 262266 02/20/09 10:06 LeRose NMR Frequency sweep
  • 262264 02/20/09 10:00 A. Camsonne Compton test Edet run 15888
  • 262261 02/20/09 09:57 V. Sulkosky DT Glitches Still Present
  • 262260 02/20/09 09:55 A. Camsonne Compton test Edet run 15888
  • 262259 02/20/09 09:50 A. Camsonne Compton test Edet run 15887
  • 262258 02/20/09 09:46 A. Camsonne Compton test Edet run 15886
  • 262255 02/20/09 09:40 Nuruzzaman Bigbite configuration changed
  • 262254 02/20/09 09:39 LeRose Target to 3He/Reference Cell vented and evacuated
  • 262253 02/20/09 09:38 Nuruzzaman Changed target,prescale factors
  • 262246 02/20/09 08:52 LeRose Target changed Reference Cell
  • 262243 02/20/09 08:08 Nuruzzaman Compton & Happex Daqs restarted
  • 262241 02/20/09 07:59 H. Gao Owl shift summary
  • 262236 02/20/09 06:06 R. Michaels end-run for 1621 actually
  • 262234 02/20/09 05:58 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 262232 02/20/09 05:34 R. Michaels absorption spectrum
  • 262223 02/20/09 04:22 R. Michaels re: Compton CMUAccum
  • 262220 02/20/09 04:16 R. Michaels NMR result (63.65%)
  • 262217 02/20/09 03:43 R. Michaels restarted Compton FADC DAQ
  • 262212 02/20/09 02:36 H. Gao Spot check (run 20282)
  • 262211 02/20/09 02:16 Moffit Restarted HAPPEX, Compton, ComptonAccum
  • 262210 02/20/09 02:09 Moffit Set golden runs to 1619 and 20282 for 10uA running.
  • 262209 02/20/09 01:22 R. Michaels BB deadtime fluctuations
  • 262208 02/20/09 01:18 R. Michaels BB rate fluctuations
  • 262207 02/20/09 00:56 R. Michaels filesplitting was off
  • 262202 02/20/09 00:39 R. Michaels re: run 1597 glitches
  • 262201 02/20/09 00:11 E.Chudakov Shift Summary
  • 262197 02/20/09 00:08 Yi Qiang Lucy Richardson Deconvolution
  • Thursday

  • 262193 02/19/09 23:54 E.Chudakov Tools
  • 262190 02/19/09 23:47 Bo Zhao NMR Frequency sweep
  • 262189 02/19/09 23:38 E.Chudakov Spot size
  • 262184 02/19/09 23:02 Moffit Restarted HAPPEX, Compton, ComptonAccum
  • 262183 02/19/09 22:55 Moffit Bigbite CODA: run with bb_buf_FADC tonight
  • 262182 02/19/09 22:52 Moffit Recompiled event_list_ppc used by ROC 5,8,9,12
  • 262181 02/19/09 22:49 E.Chudakov Moller results
  • 262174 02/19/09 20:55 adaq target pol. and Holding field current
  • 262172 02/19/09 20:54 adaq target Oven and cell temperature
  • 262150 02/19/09 20:03 O. Glamazdin BB wire chambers HV OFF
  • 262147 02/19/09 19:21 D. Parno Figure for J kinematic good electron rate
  • 262146 02/19/09 19:17 Bo Zhao Laser Absorption Spectrum
  • 262145 02/19/09 19:15 D. Parno Good electron rate for J kinematic
  • 262144 02/19/09 19:11 Bo Zhao NMR Frequency sweep
  • 262139 02/19/09 18:24 D. Parno BB golden run set to 1604
  • 262137 02/19/09 18:10 Jin Huang Re: Analyzer problem: solusion
  • 262134 02/19/09 17:52 D. Parno HAPPEX/Compton DAQs restarted
  • 262133 02/19/09 17:41 Moffit A bit more analysis on the discontinuity...
  • 262132 02/19/09 17:21 Kalyan Re: Displacement in BB Total Sum ped_sub ADCs
  • 262131 02/19/09 17:13 D. Parno Sync Check 2 discontinuity appears again
  • 262126 02/19/09 16:15 ge jin routine MWDC check
  • 262125 02/19/09 16:11 A. Camsonne Edet run 15852
  • 262123 02/19/09 16:06 D. Parno Displacement in BB Total Sum ped_sub ADCs
  • 262120 02/19/09 16:03 Saha Shift Summary
  • 262119 02/19/09 16:02 D. Parno MPS disagreement in Sync Check 1 for Run 1602
  • 262118 02/19/09 16:01 A. Camsonne Compton Edet test run 15850
  • 262117 02/19/09 15:30 D. Parno HAPPEX/Compton DAQs restarted
  • 262114 02/19/09 15:15 D. Parno BB trigger supervisor rebooted after Run 1601
  • 262113 02/19/09 15:11 mcanan 2nd_NMR_02_19_day_shift
  • 262110 02/19/09 14:54 D. Parno Helicity disagreement in Sync Check1 for Run 1601
  • 262109 02/19/09 14:36 D. Parno Sync check discontinuity still present
  • 262104 02/19/09 13:44 Saha Hall A Tools
  • 262103 02/19/09 13:43 brads re: Sync Checks not quite right for Run 1597
  • 262102 02/19/09 13:38 Saha Run 1599 SPOT++
  • 262101 02/19/09 13:31 mcanan Laser_Absorption_Spectrum_02_19_2009
  • 262100 02/19/09 13:30 mcanan Target Charts screen shot
  • 262098 02/19/09 13:26 brads Raised BB T6 threshold to -133mV (was -120mV)
  • 262096 02/19/09 12:49 ellie Sync Checks not quite right for Run 1597
  • 262093 02/19/09 12:39 ellie Runs 20248-20254 and 1564-1575 from 2/18 Swing Shift
  • 262082 02/19/09 12:23 ellie Runs 20248-20254 and 1564-1575 from 2/18 Swing Shift
  • 262081 02/19/09 11:54 ellie Left Arm VDC Efficiency Plots
  • 262078 02/19/09 10:58 mcanan 1st_NMR_02_19_day_shift
  • 262073 02/19/09 10:47 Ellie Golden run set to 20269
  • 262072 02/19/09 10:45 Moffit Analyzer problem?
  • 262071 02/19/09 10:34 brads re: Sync Check2(straight lines) giving small lines (fixed)
  • 262066 02/19/09 09:33 adaq Sync Check2(straight lines) giving small lines
  • 262065 02/19/09 09:23 adaq Re: BB cer ADCs look a bit different for first 50k events of run 1594
  • 262064 02/19/09 09:13 ellie LHRS Golden Run Set to 20268
  • 262061 02/19/09 09:05 mcnulty BB cer ADCs look a bit different for first 50k events of run 1594
  • 262056 02/19/09 08:05 Gao owl shift summmary
  • 262050 02/19/09 06:41 zhangyi NMR Frequency sweep
  • 262047 02/19/09 06:40 zhangyi NMR Frequency sweep
  • 262044 02/19/09 06:14 zhangyi front room gas pannel low prressure alarm
  • 262039 02/19/09 05:26 zhangyi halla main tools
  • 262018 02/19/09 02:39 zhangyi NMR Frequency sweep during beam trip
  • 262013 02/19/09 01:34 zhangyi beam spot
  • 262008 02/19/09 00:51 A. Camsonne Polar off on edet
  • 262007 02/19/09 00:40 A. Camsonne 2 plane only
  • 262002 02/19/09 00:00 R. Subedi Shift Summary (Swing)
  • Wednesday

  • 262000 02/18/09 23:54 A. Camsonne Run Edet in test mode 15840
  • 261999 02/18/09 23:41 R. Subedi spot++ for checklist
  • 261998 02/18/09 23:40 A. Camsonne Edet dark current
  • 261997 02/18/09 23:39 adaq L-arm VDC efficiency
  • 261996 02/18/09 23:31 R. Subedi Hall A tools
  • 261993 02/18/09 23:25 Jin Huang Re :NMR screen (alarm is cleared)
  • 261992 02/18/09 23:25 Bo Zhao log rates of different raster size
  • 261991 02/18/09 23:24 A. Camsonne Edet run 15838
  • 261990 02/18/09 23:20 R. Subedi Spot for 3x4 raster
  • 261988 02/18/09 23:11 A. Deur screenshot of target temperatures
  • 261987 02/18/09 23:09 A. Deur NMR screen (alarm for NMR sweep done won't clear)
  • 261986 02/18/09 23:08 A. Camsonne Electron detector test run 15836
  • 261985 02/18/09 23:02 R. Subedi Cycled Q2/3 (late entry)
  • 261984 02/18/09 23:00 A. Camsonne Happex and Compton FADC restarted
  • 261982 02/18/09 22:54 R. Subedi spot for raster 2x2
  • 261980 02/18/09 22:47 A. Deur He3 cell in position, NMR done (63.3%)
  • 261979 02/18/09 22:47 R. Subedi Momentum set to 1.42 GeV/c for 3He production
  • 261971 02/18/09 22:07 R. Subedi Set momentum as 0.6 GeV/c
  • 261963 02/18/09 21:37 V. Sulkosky BB Synch Check Plots
  • 261961 02/18/09 21:31 R. Subedi Set momentum to 0.7 GeV/c
  • 261957 02/18/09 21:21 A. Deur N2 reduced to 100 psig
  • 261954 02/18/09 21:10 A. Deur N2 pressure
  • 261953 02/18/09 21:03 Bo Zhao a little shift on some BB Cer ADCs
  • 261952 02/18/09 20:52 V. Sulkosky Runs 20448-20254 and 1564-1575 H2 not N2
  • 261949 02/18/09 20:43 brads BPMA Y position off by 1mm for nitrogen cell runs
  • 261948 02/18/09 20:42 A. Deur "Nitrogen" runs were actually Hydrogen runs.
  • 261947 02/18/09 20:37 brads Beam energy shifted by +1 MeV.
  • 261938 02/18/09 20:22 R. Subedi Set momentum to 1.42 GeV/c
  • 261937 02/18/09 20:03 D. Parno syncroc12 glitch reappears in Run 1383
  • 261936 02/18/09 19:56 A. Deur N2 Ref. cell vented
  • 261931 02/18/09 19:35 D. Parno Glitch in syncroc12 for Run 1386
  • 261928 02/18/09 19:23 R. Subedi Momentum set to 0.6 GeV/c for N2 run
  • 261925 02/18/09 18:58 D. Parno HAPPEX/Compton restarted
  • 261924 02/18/09 18:54 xqian new Threshold plot
  • 261923 02/18/09 18:53 Yi Qiang The pressure increasing is mainly due to the beam heating up the cell
  • 261922 02/18/09 18:52 R. Subedi spot check for N2 production
  • 261919 02/18/09 18:42 R. Subedi Momentum 0.7 GeV/c setting
  • 261918 02/18/09 18:37 R. Subedi Slight change in Run Plan
  • 261917 02/18/09 18:33 A. Deur Target back to ref position
  • 261916 02/18/09 18:25 A. Deur Target in "empty" position
  • 261913 02/18/09 18:17 Kalyan & xqi dead time study
  • 261912 02/18/09 17:52 A. Deur N2 pressure has increased slightly
  • 261910 02/18/09 17:50 D. Parno Runs 1564-1569 are with N2 reference cell target
  • 261886 02/18/09 17:28 Bo Zhao log rates of different raster size
  • 261883 02/18/09 17:14 Yi Qiang BigBite Gas Cherenkov Timing Check
  • 261881 02/18/09 17:04 A. Deur Ref cell filled with N2
  • 261880 02/18/09 17:00 R. Subedi Spot for 2x2 raster check
  • 261877 02/18/09 16:42 R. Subedi Beam Permit
  • 261876 02/18/09 16:23 Ole Hansen Analyzer updated - support for FADC
  • 261875 02/18/09 16:17 D. Parno HAPPEX/Compton DAQs restarted
  • 261873 02/18/09 15:58 Higinbotham Shift Summary
  • 261872 02/18/09 15:55 mcnulty target screen shots
  • 261869 02/18/09 15:47 Jin Huang Replay Scripts Fixed
  • 261865 02/18/09 15:36 Moffit last scaler read fixed for FADC configs.
  • 261861 02/18/09 15:20 V. Sulkosky Deadtime Check for Longitudinal Production Runs
  • 261860 02/18/09 15:19 Moffit scaler readings bogus in CODA configs that include FADC
  • 261857 02/18/09 15:15 Moffit Errors messages from the Event Builder.
  • 261854 02/18/09 15:09 Moffit Coda errors when fadc in buffered mode.
  • 261852 02/18/09 15:06 Higinbotham Controlled Access to Fix Quad
  • 261847 02/18/09 14:14 D. Parno Current Kinematic Set to J270 for Database
  • 261846 02/18/09 14:03 Yi Qiang Target Spin Rotated to 270 degrees
  • 261845 02/18/09 13:54 Higinbotham Left HRS Set To 1.43 GeV/c
  • 261844 02/18/09 13:50 ge jin routine check for MWDC
  • 261843 02/18/09 13:40 Higinbotham WARNING: Rotation Started While CODA Running
  • 261839 02/18/09 13:03 mcnulty NMR Specrta
  • 261823 02/18/09 11:23 golge Run #1549 High Dead time
  • 261822 02/18/09 11:19 golge Run # 1549 High Dead Time Junk Run
  • 261817 02/18/09 10:50 Higinbotham Hall A Tools Screen
  • 261816 02/18/09 10:49 Higinbotham Raster Size Check
  • 261813 02/18/09 10:05 V. Sulkosky Re: Restarted Compton and HAPPEX DAQs
  • 261810 02/18/09 09:47 gjin restart Compton CODA b/c flate rate
  • 261805 02/18/09 08:47 mcnulty NMR Specrta
  • 261802 02/18/09 08:43 ge jin missing golden run background in BB analysis
  • 261797 02/18/09 08:04 Ibrahim Owl Shift Summary (Feb. 18, 2009)
  • 261795 02/18/09 07:28 Nelyubin absorption spectrum
  • 261792 02/18/09 07:23 Nelyubin Oven temperature
  • 261791 02/18/09 06:58 Yi Qiang The changed is due to slightly increase in the pumping cell temperature
  • 261787 02/18/09 06:52 Nelyubin Oven temperature changing
  • 261783 02/18/09 06:33 Nelyubin absorption spectrum is changed
  • 261782 02/18/09 06:24 Nelyubin Laser Absorption Spectrum
  • 261781 02/18/09 06:24 Ibrahim Restarted the Compton and HAPPEX DAQs
  • 261780 02/18/09 06:12 Ibrahim Half Wave Plate is OUT
  • 261761 02/18/09 03:44 D. Parno BB golden run set to 1533
  • 261760 02/18/09 03:33 Nelyubin Target Holding Field and target polarization
  • 261759 02/18/09 03:31 Nelyubin Target Cell and Oven Temperature
  • 261758 02/18/09 03:23 Nelyubin NMR frequency Sweep (59.58%)
  • 261751 02/18/09 02:42 Ibrahim Half wave plate is IN
  • 261750 02/18/09 02:38 D. Parno HAPPEX/Compton DAQs restarted
  • 261747 02/18/09 01:51 D. Parno Analyzer segfaults while doing physics replay of left HRS runs
  • 261744 02/18/09 01:10 xqian Trigger problem, comments
  • 261743 02/18/09 01:03 Kalyan Preshower sum plots
  • 261742 02/18/09 01:01 xqian Trigger V
  • 261741 02/18/09 01:00 xqian Trigger problem IV
  • 261740 02/18/09 00:57 xqian Trigger problem III
  • 261739 02/18/09 00:55 xqian Trigger problem II
  • 261737 02/18/09 00:53 xqian Observation of trigger problem I
  • 261734 02/18/09 00:13 brads/kalyan Controlled access activity (late entry)
  • 261732 02/18/09 00:03 brads re: BB pmt 1 and 2 fixed (late entry)
  • 261730 02/18/09 00:00 xiaohui shift summary
  • Tuesday

  • 261727 02/17/09 23:17 deng Target moved to He-3
  • 261724 02/17/09 23:02 deng target moved to optics
  • 261723 02/17/09 22:51 deng screenshot for NMR
  • 261722 02/17/09 21:19 Moffit BB pmt 1 and 2..
  • 261720 02/17/09 21:12 deng reference cell pressure display and offset
  • 261718 02/17/09 20:50 D. Parno Missing BB Cerenkov ADCs
  • 261715 02/17/09 20:23 deng target moved to 3He production
  • 261704 02/17/09 20:07 dwang calibration run
  • 261697 02/17/09 19:27 deng target moved to reference cell
  • 261696 02/17/09 19:25 brads re: 15ns shift of BB Cerenkov group timing at the sh-grp AND for T2
  • 261693 02/17/09 19:24 dwang Raster check 1.5x2.0
  • 261692 02/17/09 19:17 V. Sulkosky Reference Pressure Gauge Readings
  • 261689 02/17/09 19:02 Moffit replaced HV crates portserver (bbps1)
  • 261688 02/17/09 18:31 debg NMR Scan
  • 261685 02/17/09 18:18 A. Camsonne Compton rates OK for production
  • 261682 02/17/09 18:15 deng target moved to He3 cell
  • 261680 02/17/09 18:12 dwang Raster run 5.5x6
  • 261676 02/17/09 18:03 dwang Raster run 6x6
  • 261671 02/17/09 17:54 dwang Raster size check 2 x 2
  • 261665 02/17/09 17:32 xqian Angular dependence in the energy
  • 261664 02/17/09 17:14 deng target moved to BeO
  • 261663 02/17/09 17:09 xqian Energy & Trigger energy vs x position of shower counter
  • 261662 02/17/09 17:09 xiaohui beam chart during BCM calibration
  • 261661 02/17/09 17:02 xiaohui BCM calibration run 30uA
  • 261656 02/17/09 16:50 xiaohui BCM calibration run
  • 261653 02/17/09 16:17 xiaohui HV turned off near the end of the swing shift last night
  • 261649 02/17/09 16:07 X. Zheng Shift Summary
  • 261648 02/17/09 16:00 A. Camsonne Compton Run15799 standalone polling electron
  • 261647 02/17/09 15:33 Moffit New CODA configs for ROC16 readout
  • 261646 02/17/09 15:24 Shahinyan Target
  • 261640 02/17/09 15:00 X. Zheng MCC update again
  • 261639 02/17/09 14:43 X. Zheng MCC update
  • 261638 02/17/09 14:43 deng TransPvdis run 30813
  • 261637 02/17/09 14:38 X. Zheng BB shower threshold T6 set to -120mV
  • 261636 02/17/09 13:55 Yi Qiang One flaky cable of down stream pick up coils, but not solving the problem
  • 261635 02/17/09 13:53 X. Zheng Going back to beam permit
  • 261632 02/17/09 13:36 Higinbotham BigBite Chambers Off
  • 261630 02/17/09 13:25 JinGe routine MWDC check
  • 261628 02/17/09 13:12 Yi Qiang EPR FM verification, starting polarization: 45.71%
  • 261626 02/17/09 12:58 Yi Qiang Target Spin Rotated to 0 degree
  • 261625 02/17/09 12:47 V. Sulkosky Experiment Summary for Feb 16, Owl - Feb 17, Owl
  • 261624 02/17/09 12:43 Yi Qiang NMR Cross Calibration for transverse pumping
  • 261623 02/17/09 12:28 brads BB cerenkov PMTs 1--8, 11--18 split to both 1881 and FADC
  • 261614 02/17/09 12:20 brads 20ns added to BB Cerenkov group timing at the sh-grp AND for T2
  • 261612 02/17/09 12:16 Yi Qiang COMET laser power test
  • 261609 02/17/09 11:36 Yi Qiang Internal Temperature Test for Transverse Pumping
  • 261606 02/17/09 10:11 Yi Qiang EPR Calibration with Transverse Pumping
  • 261605 02/17/09 09:05 Yi Qiang Oven temperature set to 240C
  • 261604 02/17/09 08:59 V. Sulkosky Cosmic Runs 1510-1512 Part of MWDC Off
  • 261603 02/17/09 08:51 V. Sulkosky MWDC HV's Were Off
  • 261600 02/17/09 08:09 bogdanw Next step for GC analysis
  • 261599 02/17/09 08:01 D. Parno Owl shift summary 02/17
  • 261592 02/17/09 06:12 D. Parno Magnet problems due to bad power supply
  • 261591 02/17/09 05:35 Yi Qiang More results of BB Gas Chrenkov analysis from run 1479-1482
  • 261582 02/17/09 03:26 D. Parno Magnet work ongoing
  • 261581 02/17/09 02:17 D. Parno Controlled access for magnet problem
  • 261580 02/17/09 01:05 ycwang MCC is shooting up a beam poblem
  • 261579 02/17/09 00:10 Xiaohui Shift Summary
  • Monday

  • 261577 02/16/09 23:32 xiaohui target moved to empty
  • 261576 02/16/09 23:29 Kalyan new histograms added to online replay
  • 261575 02/16/09 23:15 xiaohui beam not stable
  • 261574 02/16/09 22:44 xiaohui beam chart
  • 261573 02/16/09 22:40 xiaohui HV trip
  • 261568 02/16/09 21:52 deng Screenshot for NMR
  • 261565 02/16/09 21:30 deng reference cell is vacum
  • 261564 02/16/09 21:30 bogdanw preshower path in electronic
  • 261563 02/16/09 21:20 brads re: HW vs. SW row sums for run 1482 (after gain re-calibration)
  • 261562 02/16/09 20:59 Jin Huang BB Physics Replay Ready
  • 261559 02/16/09 20:32 xqian Threshold Scan of the BigBite T6 Trigger
  • 261546 02/16/09 19:48 xqian & Kaly The Software Sum should be PRESHOWER + SHOWER
  • 261545 02/16/09 19:13 Kalyan Run 1502-1505 are BB theshold scan runs
  • 261532 02/16/09 18:47 D. Parno LHRS electron rates as expected for L90 kinematic
  • 261530 02/16/09 18:44 xqian Threshold on energy now
  • 261528 02/16/09 18:00 Yi Qiang EPR FM sweep verified the success of the spin rotation
  • 261527 02/16/09 17:52 Yi Qiang Target spin direction rotated to 270 degrees: Transverse -, 61.22%
  • 261522 02/16/09 17:05 D. Parno BB Cerenkov ADC shift in Run 1496
  • 261521 02/16/09 16:48 deng screenshot for NMR
  • 261513 02/16/09 15:41 Yi Qiang The Runs used in the analysis are 1406-1433
  • 261512 02/16/09 15:34 D. Parno LHRS golden run is 20217
  • 261511 02/16/09 15:30 brads re: Analysis of Bigbite Gas Cherenkov
  • 261510 02/16/09 15:24 A. Camsonne Compton rates ok
  • 261509 02/16/09 15:20 Kalyan Pedestals for BB updated
  • 261508 02/16/09 14:57 D. Parno BB Deadtime plots exist now
  • 261503 02/16/09 14:40 A. Camsonne Compton background increased a bit
  • 261500 02/16/09 14:22 deng Stopped Transpvdis # 30807
  • 261497 02/16/09 14:14 golge BB 1493 dead time plot
  • 261496 02/16/09 14:11 R. Michaels SPOT++ check
  • 261494 02/16/09 13:55 bogdanw Low beam current => pedestals are narrow
  • 261492 02/16/09 13:48 R. Michaels run with TransPvdis
  • 261491 02/16/09 13:45 Kalyan Wiki entry for shower calibration with cosmics
  • 261488 02/16/09 12:57 ellie Re: large change in BB ADCs
  • 261487 02/16/09 12:22 R. Michaels BB HV trip / reset
  • 261478 02/16/09 12:15 R. Michaels large change in BB ADCs
  • 261475 02/16/09 12:00 jpchen back to 15 uA production run
  • 261472 02/16/09 11:48 jpchen 1 uA for Bogdan's test
  • 261469 02/16/09 11:40 R. Michaels NMR day (61.61%)
  • 261466 02/16/09 11:08 ellie Left Arm p not changed in Start Run (runs 1485-1490)
  • 261465 02/16/09 11:00 Moffit adjusted beam energy in db_run.dat for a few runs
  • 261464 02/16/09 10:39 Kalyan New database directory created 20090216
  • 261459 02/16/09 09:22 jpchen MCC do orbit check
  • 261456 02/16/09 09:15 Moffit xscaler rates for 15uA, Kinematic L
  • 261455 02/16/09 08:49 jpchen HV trip, L14.6 (chamber)
  • 261450 02/16/09 08:03 Kumbartzki Owl shift summary
  • 261448 02/16/09 07:55 R. Michaels scalers for run 20215
  • 261447 02/16/09 07:46 Kumbartzki too many BB-HV alarms
  • 261442 02/16/09 06:03 kumbartzki raster 20217
  • 261439 02/16/09 05:49 yawei Target: NMR for Trans. 3He pol. = 61.34% at 05:47 AM Feb.16, 2009
  • 261424 02/16/09 04:01 brads HW vs. SW row sums for run 1482 (after gain re-calibration)
  • 261417 02/16/09 02:21 D. Parno BB golden run set to 1479
  • 261414 02/16/09 02:05 xqian Some naive thinkings
  • 261408 02/16/09 01:16 yawei Target Laser absorption spectrum
  • 261407 02/16/09 01:15 yawei Target: NMR for Trans. 3He pol. = 62.17% at 01:09 AM Feb.16, 2009
  • 261405 02/16/09 01:14 Moffit Negative u10 and d10 coulombs from scaler roll over..
  • 261404 02/16/09 01:10 Muangma raster sport check run 20214 size 4*4
  • 261402 02/16/09 01:06 brads BB T6 threshold set to -120mV
  • 261400 02/16/09 01:04 Muangma raster spot check LHRS 20213
  • 261398 02/16/09 01:03 D. Parno Run 1471: Cerenkov TDCs vanish for both T2 and T6 events
  • 261395 02/16/09 00:54 Kalyan/Yawei BigBite HV updated based on the cosmic calibration run 1476
  • 261391 02/16/09 00:34 muangma spot run LHRS 20210
  • 261389 02/16/09 00:28 Yi Qiang Analysis of Bigbite Gas Cherenkov
  • 261387 02/16/09 00:27 Jin Huang Correction: Software sum of total sum 19
  • 261386 02/16/09 00:24 yawei move target to optics
  • 261385 02/16/09 00:23 brads HW vs. SW sums for run 1444 (before Saturday's access)
  • 261384 02/16/09 00:18 kumbartzki Harp B
  • 261383 02/16/09 00:16 brads HW vs. SW sums for run 1470 (before Sunday's access)
  • 261382 02/16/09 00:12 kumbartzki Harpscan
  • 261380 02/16/09 00:01 xiaohui Shift Summary
  • Sunday

  • 261377 02/15/09 23:20 D. Parno Cerenkov TDCs vanish for high T6 threshold
  • 261376 02/15/09 23:14 xiaohui another hour for cosmic test
  • 261375 02/15/09 23:09 deng target moved to empty position
  • 261374 02/15/09 23:08 xiaohui hall A tool
  • 261373 02/15/09 23:08 D. Parno BB golden run set to 1470
  • 261371 02/15/09 23:06 Jin Huang Software sum of total sum 19 is OK
  • 261369 02/15/09 23:01 brads HV settings for new PMT 12 in BB Cerenkov
  • 261367 02/15/09 22:42 brads BB Cerenkov base + PMT on #12 replaced
  • 261364 02/15/09 22:06 xqian Be careful with interpreting these plots
  • 261359 02/15/09 21:41 Yi Qiang Oven temperature reduced to 235C
  • 261357 02/15/09 21:15 D. Parno Wiring additions to Compton FADC DAQ
  • 261348 02/15/09 20:06 xiaohui control access
  • 261345 02/15/09 19:51 Kalyan Run 1471 is a special run with T6 threshold set to -250mV
  • 261340 02/15/09 19:31 deng screenshot for NMR
  • 261339 02/15/09 19:25 posik production runs (1449->1464)
  • 261336 02/15/09 19:07 A. Camsonne Compton central crystal HV current correlation with rate
  • 261331 02/15/09 18:08 Moffit End of run for u10 and d10 ungated is negative.
  • 261328 02/15/09 17:26 dwang BB run 1466 hits lower
  • 261322 02/15/09 15:59 Kumbartzki Day shift summary
  • 261319 02/15/09 15:19 Yunxin Ye Abs_spectr_021509_day
  • 261316 02/15/09 14:54 Yunxin Ye NMR Measurement
  • 261313 02/15/09 14:49 Zhihong Ye Reboot Coda and ROC8
  • 261309 02/15/09 14:04 kumbartzki BB HV trip
  • 261304 02/15/09 13:51 posik Energy deposited in BB shower
  • 261303 02/15/09 13:33 Moffit Added BB scaler replay to shiftworker replay.
  • 261302 02/15/09 13:30 kumbartzki Hall tools
  • 261301 02/15/09 13:27 kumbartzki spot check
  • 261297 02/15/09 12:47 Moffit Trouble with Compton DAQ
  • 261283 02/15/09 10:22 Zhihong Ye Did not stup run 20205 while NMR measurement is taking.
  • 261281 02/15/09 10:21 Yunxin Ye NMR measurement at 10:13 am
  • 261277 02/15/09 09:58 posik BB cer TDC change
  • 261274 02/15/09 09:17 kumbartzki silent alarms hv_giu continue
  • 261271 02/15/09 08:36 Jin Huang No alarm sound for BB HV GUI
  • 261270 02/15/09 08:27 Jin Huang Shift Summary Feb15 Owl
  • 261267 02/15/09 08:24 Jin Huang Access Summary: Problem with BB Trigger
  • 261265 02/15/09 07:54 Jin Huang BB Shower hit after NIM bin power fix
  • 261262 02/15/09 07:25 Jin Huang HV GUI Alarmed again
  • 261260 02/15/09 07:10 Jin Huang BB Threshold scan
  • 261259 02/15/09 07:07 Jin Huang No more alarm for BB HV GUI lost connection
  • 261256 02/15/09 06:30 Jin Huang Spot check: Clear
  • 261253 02/15/09 06:21 yawei Target Laser absorption spectrum
  • 261251 02/15/09 06:19 yawei Target: NMR for Trans. 3He pol. = 63.66% at 06:17 AM Feb.15, 2009
  • 261250 02/15/09 06:16 yawei target moved to 3He
  • 261249 02/15/09 06:15 Jin Huang Beam position and raster checking with BeO camera
  • 261248 02/15/09 06:12 Jin Huang Harp scan result
  • 261246 02/15/09 06:05 Jin Huang Target -> Optics
  • 261244 02/15/09 05:53 Jin Huang Access Over; Starting Harp Scan
  • 261226 02/15/09 00:00 R. Subedi Shift Summary (Swing)
  • Saturday

  • 261225 02/14/09 23:24 R. Subedi No beam for a while
  • 261224 02/14/09 23:18 A. Rakhman NMR frequency sweep screen
  • 261223 02/14/09 23:17 A. Rakhman Laser Absorption Spectrum
  • 261222 02/14/09 23:14 A. Rakhman Target screen strip charts
  • 261221 02/14/09 23:12 A. Rakhman Target Status
  • 261220 02/14/09 22:45 brads Pedestal subtracted EDTM pulser signal in BB shower row-sums
  • 261215 02/14/09 22:13 Moffit Frequent connections lost to HV portserver
  • 261208 02/14/09 20:32 R. Subedi spot++
  • 261207 02/14/09 20:28 R. Subedi Spot++
  • 261198 02/14/09 19:58 A.Rakhman Laser Absorption Spectrum
  • 261197 02/14/09 19:55 A.Rakhman NMR screenshot
  • 261190 02/14/09 19:17 Jack (techs) Right Arm
  • 261187 02/14/09 18:52 Moffit Changed golden for BB to 1405
  • 261186 02/14/09 18:32 Moffit Restarted HVgui
  • 261185 02/14/09 18:21 R. Subedi BB HV trip
  • 261182 02/14/09 18:12 D. Parno Compton rates very good
  • 261181 02/14/09 17:44 Moffit Wrong energy readback for start of 20197
  • 261178 02/14/09 17:14 D. Parno LHRS physics golden run now set to 20197
  • 261177 02/14/09 17:13 A.Rakhman NMR screenshot
  • 261176 02/14/09 17:00 A.Rakhman Target Screen capture
  • 261174 02/14/09 16:58 A.Rakhman Laser Absorption Spectrum
  • 261173 02/14/09 16:57 D. Parno LHRS Golden Run = 20197
  • 261168 02/14/09 16:21 Kalyan T6 threshold is set back to -80mV
  • 261164 02/14/09 16:03 Kumbartzki Day shift summary
  • 261163 02/14/09 15:52 Kalyan Run 1435 is a special run with T6 threshold set to 215mV --> 800MeV
  • 261160 02/14/09 15:26 posik absorption spectrum
  • 261158 02/14/09 15:22 posik NMR polarization 61.74%
  • 261154 02/14/09 15:01 kumbartzki Hall A tool screen
  • 261153 02/14/09 14:58 kumbartzki spot run 20195
  • 261126 02/14/09 10:42 posik NMR polarization 61.19
  • 261110 02/14/09 08:01 Jin Huang Shift Summary for Feb14 Owl
  • 261109 02/14/09 07:50 Jin Huang Target Spin Flip X- -> X+
  • 261088 02/14/09 05:07 Jin Huang Hall A Tools, Feb 14 Owl
  • 261087 02/14/09 05:06 Jin Huang Spot++ Feb14 Owl
  • 261084 02/14/09 04:55 D. Parno Compton vertical tune done
  • 261083 02/14/09 04:54 yawei target moved to 3He
  • 261082 02/14/09 04:38 D. Parno Card 11 Slot 1 is the Compton
  • 261081 02/14/09 04:37 Jin Huang Target out of beam for Compton tuning
  • 261078 02/14/09 04:30 D. Parno Deposited-energy gap in BB shower rows
  • 261067 02/14/09 04:03 brads The intermittant beamline HV alarm is due to card 11, chan 0 (Compton photon detector?)
  • 261066 02/14/09 03:55 yawei Absorption line spectrum
  • 261065 02/14/09 03:52 D. Parno BB Acceptance with Trigger 6
  • 261064 02/14/09 03:47 yawei Polarization is 60.43%
  • 261063 02/14/09 03:46 yawei Target: NMR for Trans. 3He pol. = 55.13% at 03:40 AM Feb.14, 2009
  • 261058 02/14/09 03:29 Jin Huang Dead time after thr change / new scalar replay
  • 261055 02/14/09 03:05 Jin Huang Multiple BB HV Gui problem
  • 261054 02/14/09 02:59 Kalyan BB acceptance issues
  • 261050 02/14/09 02:22 Jin Huang BB HV Trip
  • 261046 02/14/09 01:56 D. Parno BB golden run set to 1403
  • 261035 02/14/09 00:53 Moffit Added SpinInfo table to d2n database... other changes.
  • 261034 02/14/09 00:47 Moffit Removed T1 from BigBite datamon
  • 261030 02/14/09 00:24 brads T6 threshold lowered to -80mV
  • 261028 02/14/09 00:12 D. Parno Set golden runs to 20180, 1402
  • 261027 02/14/09 00:06 V. Sulkosky Left HRS Physics Replay Ready
  • 261026 02/14/09 00:02 D. Parno LHRS Cerenkov changes at new momentum setting
  • Friday

  • 261024 02/13/09 23:57 Kalyan Run normalization and offline deadtime set up L-HRS
  • 261023 02/13/09 23:56 Kumbartzki Swing shift summary
  • 261022 02/13/09 23:42 xqian forget plot
  • 261021 02/13/09 23:41 xqian Clear Hole in acceptance for T6 trigger
  • 261020 02/13/09 23:31 kumbartzki spot after restart
  • 261019 02/13/09 23:27 mcanan 3rd_NMR_02_13_swing_shift
  • 261016 02/13/09 23:23 mcanan 3rd_NMR_02_13_swing_shift
  • 261013 02/13/09 23:14 mcanan Target Motion
  • 261010 02/13/09 23:02 mcanan Target Motion
  • 261009 02/13/09 23:02 brads Controlled access to shift BB retimed L1A by -25ns
  • 261006 02/13/09 22:10 mcanan Laser Absorption Spectrum 02_13_2009
  • 261005 02/13/09 22:08 mcanan target oven temperature
  • 261004 02/13/09 22:07 mcanan holding field current
  • 261003 02/13/09 22:06 mcanan target_cell_temperature
  • 261002 02/13/09 22:05 brads BB Cerenkov disc. group signal was 50ns, now 30ns
  • 261001 02/13/09 22:00 kumbartzki final spot after break
  • 261000 02/13/09 21:59 mcanan 2nd_NMR_02_13_swing_shift (screen shot)
  • 260999 02/13/09 21:58 mcanan 2nd_NMR_02_13_swing_shift
  • 260988 02/13/09 21:30 mcanan Target Motion
  • 260987 02/13/09 21:26 kumbartzki raster check after Access
  • 260984 02/13/09 21:19 mcanan Target motion
  • 260983 02/13/09 21:15 brads BB Cerenkov signals moved to 1881 ADC, ROC8, slot 9, ch 32--63
  • 260982 02/13/09 21:14 D. Parno Gain Shift in Compton Photon Detector
  • 260981 02/13/09 21:11 brads 60Hz ADC filter for BB Cer PMT 6 was bad, replaced filter
  • 260979 02/13/09 21:02 Kalyan Re: Controlled access activity - DAQ changes
  • 260973 02/13/09 20:47 V. Sulkosky BigBite GC HV Adjusted
  • 260969 02/13/09 20:40 Kalyan Correction: Re:Controlled access activity - DAQ changes
  • 260968 02/13/09 20:33 Kalyan Controlled access activity - DAQ changes
  • 260945 02/13/09 19:43 kumbartzki New momentum 0.700
  • 260941 02/13/09 19:04 Yi Qiang Rutgers COMET now in use
  • 260940 02/13/09 18:31 mcanan 1st_NMR_02_13_swing_shift
  • 260937 02/13/09 18:16 xqian Naive tracking efficiency is high
  • 260932 02/13/09 17:22 kumbartzki shift checklist tools screen
  • 260931 02/13/09 17:20 kumbartzki spot at checklist time
  • 260930 02/13/09 17:12 O. Glamazdin Corrected Moller results on 11 Feb 2009
  • 260925 02/13/09 16:18 X. Zheng day shift summary
  • 260923 02/13/09 16:03 R. Michaels adjust SPLITMB for bigbite_buf
  • 260920 02/13/09 15:54 LeRose Laser Absorption Spectrum
  • 260917 02/13/09 15:13 Yi Qiang More Rutgers COMET test with new fiber
  • 260914 02/13/09 15:03 Kalyan offline deadtime is close to online DT(from datamon/end-run)
  • 260913 02/13/09 14:20 LeRose Target NMR Sweep
  • 260903 02/13/09 13:04 Kalyan Re: BB - some histograms empty
  • 260901 02/13/09 12:45 E.Chudakov BB - some histograms empty
  • 260890 02/13/09 12:05 Yi Qiang Rutgers COMET power test with UVa jumper cable
  • 260885 02/13/09 10:44 LeRose Target NMR Sweep
  • 260880 02/13/09 10:26 X. Zheng Bogdan's test (half beam current)
  • 260875 02/13/09 09:14 X. Zheng Hall A tools screen
  • 260874 02/13/09 09:12 X. Zheng spot++
  • 260868 02/13/09 07:58 Jin Huang Shift Summary for Feb13 Owl
  • 260863 02/13/09 06:38 Jin Huan No End of run for 1374. Rebooting Roc10
  • 260861 02/13/09 06:14 Jin Huang Reboots for LHRS
  • 260858 02/13/09 05:56 yawei Target: NMR for Trans. 3He pol. = 55.13% at 05:55 AM Feb.13, 2009
  • 260835 02/13/09 03:07 Jin Huang Hall A tools. Owl shift
  • 260834 02/13/09 03:07 Jin Huang Spot++ Owl
  • 260827 02/13/09 01:58 Jin Huang Run 1366 1368 30158 -> 1uA
  • 260825 02/13/09 01:46 Jin BB MWDC Trip during
  • 260824 02/13/09 01:46 Kalyan Re: BB Deadtime Discrepancy
  • 260823 02/13/09 01:45 V. Sulkosky Re: BB Deadtime Discrepancy
  • 260813 02/13/09 01:24 V. Sulkosky Started New HAPPEX Run 30778 (Late Entry)
  • 260812 02/13/09 01:21 yawei Absorption line spectrum
  • 260811 02/13/09 01:15 yawei Target: NMR for Trans. 3He pol. = 55.22% at 01:02AM Feb.13, 2009
  • 260810 02/13/09 01:12 Jin Huang Run 1364; 20157 --> 5uA
  • 260803 02/13/09 00:58 V. Sulkosky BB Deadtime Discrepancy
  • 260800 02/13/09 00:31 V. Sulkosky Golden Root File -> Run #1360 (BB)
  • 260799 02/13/09 00:28 V. Sulkosky Re: BB scint right ped_sub ADCs
  • 260798 02/13/09 00:26 brads OWL shift requests
  • 260796 02/13/09 00:00 ge jin BB scint right ped_sub ADCs
  • Thursday

  • 260795 02/12/09 23:59 kumbartzki Swing shift Summary
  • 260790 02/12/09 23:19 ge jin a wire in wire chamber abnormal
  • 260789 02/12/09 23:13 A. Camsonne Compton vertical position tune.
  • 260788 02/12/09 22:58 Yunxin Ye target moved from empty to Helium-3
  • 260785 02/12/09 22:44 Yunxin Ye Abs_spectr_021209_swing
  • 260784 02/12/09 22:38 Yunxin Ye Moving target to empty
  • 260783 02/12/09 22:35 Yunxin Ye NMR sweeping-pol. from 56.38 to 56.29
  • 260780 02/12/09 22:02 ge jin hall a tools screen
  • 260777 02/12/09 21:29 A. Camsonne Previous Compton beam parameters
  • 260776 02/12/09 21:22 Kumbartzki spot of production run 20154
  • 260775 02/12/09 21:20 kumbartzki spot 6x6
  • 260774 02/12/09 21:17 ge jin rates of a raster size
  • 260769 02/12/09 20:30 Ge jin rates of different raster sizes
  • 260766 02/12/09 20:17 kumbartzki spot sizes on He3 cell
  • 260753 02/12/09 19:55 Yunxin Ye Target moving
  • 260751 02/12/09 19:49 kumbartzki Spot
  • 260749 02/12/09 19:44 kumbartzki Spot
  • 260746 02/12/09 19:28 Yunxiu Ye Target moving
  • 260745 02/12/09 19:28 Kumbartzki harp scan
  • 260744 02/12/09 19:25 D. Parno BB raster correlations added to online plots
  • 260743 02/12/09 19:25 kumbartzki harp scan
  • 260742 02/12/09 18:48 Jin Target -> Empty
  • 260741 02/12/09 18:46 Jin Target Polarization was 56% before beam tuning
  • 260740 02/12/09 18:35 brads Hall back to Beam Permit
  • 260735 02/12/09 18:03 brads Lots of noise on BB T8 gate upstream of TS and scalers (and probably T6) due to T7 input
  • 260730 02/12/09 17:52 Moffit Added more wait time for ROC5 ADCs
  • 260728 02/12/09 17:49 Moffit Added Raster Y to ROC5
  • 260724 02/12/09 17:29 brads T8 events are now analyzed in the BigBite replays...
  • 260714 02/12/09 16:47 V. Sulkosky Re: Updated Trigger Latch Pattern Detector Map
  • 260698 02/12/09 16:17 Kalyan Updated BB scintillator HV
  • 260691 02/12/09 15:54 Higinbotham Shift Summary (day shift)
  • 260690 02/12/09 15:54 brads BB Cer PMT 17 is noiser than the rest
  • 260685 02/12/09 15:46 adaq Hall A Tools Screen
  • 260678 02/12/09 15:34 brads Snapshot of pedestal shift with different v792 iped values (Cerenkov ADC)
  • 260677 02/12/09 15:32 Higinbotham Access While No Beam
  • 260672 02/12/09 15:20 brads BB Cerenkov ADC weirdness
  • 260657 02/12/09 14:54 Moffit Changed LHRS TS FrontEnd BUSY to 60us
  • 260646 02/12/09 13:29 brads Changes to ROC5 crl code (pedestals will change)
  • 260633 02/12/09 12:01 Katich Spare d2n cell 'MOSS' prepared
  • 260632 02/12/09 11:59 Katich Spare 3He bottle has arrived
  • 260627 02/12/09 08:51 Higinbotham North Linac access.
  • 260624 02/12/09 08:09 A. Tobias Owl Shift Summary
  • 260622 02/12/09 07:52 A. Tobias spot++ LHRS run 20146
  • 260621 02/12/09 07:49 A. Tobias Increase beam current to 15uA: Prescales Set for LHRS & BB
  • 260614 02/12/09 06:10 A. Tobias Hall A Tools after iocha22 recovery
  • 260613 02/12/09 06:01 A. Tobias spot++ LHRS run 20145
  • 260608 02/12/09 05:31 V. Sulkosky Golden Root File -> Run #1310 (BB) and # 20144 (Left HRS)
  • 260607 02/12/09 05:12 A. Tobias Target: Laser Absorption Spectrum Owl Shift
  • 260606 02/12/09 05:03 V. Sulkosky Compton and Charge Asymmmetry Issues
  • 260605 02/12/09 05:02 A. Tobias spot++ for Run LHRS 20144 during iocha22 outage
  • 260604 02/12/09 04:58 yawei roc5 not respond, kcoda to bring it back, no e-log entry for the end run
  • 260602 02/12/09 04:54 A. Tobias iocha22 & iocha1 dead: Called MCC to reboot Gateway
  • 260601 02/12/09 04:40 A. Tobias Got Compton & Charge asymmetry disconnects??
  • 260590 02/12/09 04:01 A. Tobias BB: HV (S14 & S15) for Wire Chamber Turned back ON
  • 260585 02/12/09 03:49 V. Sulkosky Updated db_DL.dat Trigger Bit Pattern
  • 260584 02/12/09 03:48 A. Tobias Target: NMR for Trans. 3He pol. = 56.95% at 3:39AM Feb.12, 2009
  • 260582 02/12/09 03:36 A. Tobias spot++ on Pol. 3He w/ 6x7 MCC and 2uA
  • 260580 02/12/09 03:30 A. Tobias spot++ on Pol. 3He w/ 4x5.2 MCC and 2uA
  • 260576 02/12/09 03:24 A. Tobias spot++ on Pol3 He w/ 2x2.6MCC and 2uA
  • 260574 02/12/09 03:14 A. Tobias Target: Moved to Trans. Pol. 3He
  • 260572 02/12/09 03:09 A. Tobias spot++ on Carbon/BeO
  • 260570 02/12/09 02:59 V. Sulkosky Harp Scan Results
  • 260569 02/12/09 02:10 R. Michaels re: 1875 missing gate
  • 260567 02/12/09 00:08 Kumbartzki Swing shift summary
  • Wednesday

  • 260565 02/11/09 23:42 Yunxiu Ye Moving target to optics
  • 260564 02/11/09 23:27 Kumbartzki BPM readings
  • 260563 02/11/09 22:56 V. Sulkosky Re: ROC 4, Slot 18 (1875 TDC) Missing Gates in Buffered Mode
  • 260562 02/11/09 22:54 Yunxin Ye Laser_abs_spectr_021109_swing
  • 260561 02/11/09 22:42 Moffit ROC 5 errors from 1308
  • 260560 02/11/09 22:42 V. Sulkosky Test Runs for Left HRS Buffering
  • 260554 02/11/09 21:45 O. Glamazdin Moller results
  • 260551 02/11/09 21:23 V. Sulkosky DecData Class Split for BigBite Replay
  • 260550 02/11/09 20:54 R. Michaels re: missing synch on L-HRS
  • 260549 02/11/09 20:46 R. Michaels missing synch on L-HRS
  • 260548 02/11/09 20:32 Jin Huang Target Direction in d2n Data
  • 260537 02/11/09 18:42 Jin Huang Transversity Online File Backup/Clean: Done
  • 260536 02/11/09 17:12 O. Glamazdin BB wire chambers HV OFF
  • 260535 02/11/09 17:08 V. Sulkosky V. Sulkosky
  • 260534 02/11/09 17:03 V. Sulkosky Warning Messages in THaBPM and THaRaster
  • 260532 02/11/09 16:05 V. Sulkosky Re: spot++ fix for 2-DAQ running
  • 260531 02/11/09 16:04 Kalyan BB shower x vs y
  • 260530 02/11/09 15:57 Ole Hansen Day Shift Summary
  • 260516 02/11/09 15:18 brads MWDC thresholds set to 3.9V, 4,8V, 4.8V
  • 260509 02/11/09 13:10 Moffit CAMAC startup script executed.
  • 260500 02/11/09 12:10 Ole Hansen Beam Permit
  • 260498 02/11/09 11:50 Yi Qiang NMR from transverse pumping: 52%
  • 260496 02/11/09 11:35 Yi Qiang Bigbite current set to 710A, holding field adjusted
  • 260495 02/11/09 11:30 Ole Hansen BigBite CAMAC crate work/power cycle
  • 260494 02/11/09 11:30 Yi/Jin The rest two Q-wave plates in Longitudinal S-wave were adjusted
  • 260493 02/11/09 11:25 Ole Hansen BigBite magnet to 710 A
  • 260492 02/11/09 11:17 D. Parno Compton Counting DAQ Signal Attenuated
  • 260481 02/11/09 10:10 V. Sulkosky Re: Left HRS Triggers for Normalization Scalers
  • 260480 02/11/09 09:58 V. Sulkosky T3 and T4 Unplugged from BigBite TS
  • 260479 02/11/09 09:56 V. Sulkosky DAQ back to 2-DAQ Mode
  • 260476 02/11/09 09:36 Ole Hansen LHRS to 600 MeV/c
  • 260467 02/11/09 09:04 Yi Qiang Field setting changed back for negative Bigbite polarity
  • 260465 02/11/09 08:23 A. Tobias Target: Holding field current changed
  • 260464 02/11/09 08:20 Fansler/Ewin Polarity change
  • 260463 02/11/09 08:17 A. Tobias Target: NMR sweep for Trans. 3He not in correct frequency range?
  • 260461 02/11/09 07:49 Yoomin Owl Shift Summary
  • 260460 02/11/09 07:16 Yoomin L-HRS Q,D current to zero
  • 260457 02/11/09 06:43 Yoomin BB HV trip
  • 260455 02/11/09 06:40 Yoomin Compton dead
  • 260454 02/11/09 06:27 Navaphon beam current 4 uA
  • 260451 02/11/09 06:05 A. Tobias Target: Laser Absorption Spectrum Owl Shift
  • 260448 02/11/09 05:32 Yoomin Beam Current for run 1291=12uA
  • 260443 02/11/09 05:09 Yoomin Target Camera without raster
  • 260441 02/11/09 04:30 A. Tobias Target: Moved to Reference Cell w/ 133 psig H2 gas.
  • 260439 02/11/09 04:22 Yoomin Optics target run 1286, 1288
  • 260430 02/11/09 04:10 Yoomin BB sieve is OUT
  • 260428 02/11/09 04:00 A. Tobias Target: Moved to Optics/BeO
  • 260426 02/11/09 03:41 Yoomin Q2/3 Momentum change during the run
  • 260422 02/11/09 03:25 Yoomin To take the sieve OUT
  • 260421 02/11/09 03:21 brads BB Sieve OUT
  • 260420 02/11/09 03:16 Yoomin Controlled Access
  • 260414 02/11/09 02:17 Yoomin Beam size for BB optics study
  • 260411 02/11/09 01:56 Yoomin Starting BB optics
  • 260410 02/11/09 01:46 Yoomin spot++
  • 260408 02/11/09 01:43 A. Tobias Target: Reference Cell w/ 133 psig H2 gas.
  • 260405 02/11/09 01:36 Yoomin CODA died and rebooted
  • 260401 02/11/09 01:28 Yoomin Beam check with Optics w/BeO
  • 260399 02/11/09 01:25 Yoomin BB trip
  • 260398 02/11/09 01:23 Yoomin BCM log file
  • 260395 02/11/09 01:20 Yoomin Give up BCM calibration
  • 260393 02/11/09 00:40 Yi Qiang Q-wave plate calibration
  • 260392 02/11/09 00:34 Moffit HAPPEX DAQ: run 30772 for BCM calibration
  • 260391 02/11/09 00:25 Yoomin L-HRS Q3 keeps oscillating
  • 260390 02/11/09 00:03 zhangyi Swing shift summary
  • Tuesday

  • 260387 02/10/09 23:05 Yi Qiang Target -> Empty
  • 260381 02/10/09 22:01 brads Beam energy readbacks are working again
  • 260380 02/10/09 22:00 brads Oven temps still oscillating
  • 260377 02/10/09 21:41 brads Beam energy IOC is down, MCC notified
  • 260376 02/10/09 21:33 Diana Parno Compton Photon HV back to -1800 V
  • 260374 02/10/09 21:25 zhangyi fill H2 in Ref. cell. 133psi
  • 260373 02/10/09 21:23 Shahinyan Target move
  • 260371 02/10/09 21:06 Yi Qiang Summary of target optics/cell work since d2n
  • 260368 02/10/09 21:01 brads Oven Temps oscillating
  • 260367 02/10/09 21:00 xqian New hitting efficiency
  • 260366 02/10/09 20:58 brads BB Cerenkov T2 timing re-measured, looks good
  • 260364 02/10/09 20:44 Shahinyan target move
  • 260362 02/10/09 20:41 zhangyi rates under 2 uA beam with a 6x5 raster
  • 260359 02/10/09 20:38 zhangyi rates under 2 uA, 4x4 raster
  • 260357 02/10/09 20:35 Yi/Jin Rotatable Q-wave plates installed into Longitudinal line
  • 260353 02/10/09 20:29 zhangyi rates under 2 uA with a 2x2 raster
  • 260351 02/10/09 20:24 shahinyan Target move
  • 260350 02/10/09 20:19 brads Controlled Access finished 8pm
  • 260349 02/10/09 20:18 zhangyi spot++ on BeO
  • 260348 02/10/09 20:18 brads BB thresholds restored
  • 260337 02/10/09 20:03 Shainyan Target move
  • 260336 02/10/09 19:08 brads BigBite Sieve IN
  • 260335 02/10/09 18:15 zhangyi BB polarity change
  • 260334 02/10/09 18:03 Kalyan Total energy in BB calorimeter for H2 elastic run
  • 260333 02/10/09 17:55 brads BB T1, T6 thresholds set to -10mV for tests
  • 260331 02/10/09 17:13 Kalyan Hydrogen elastic run 1258 to 1262
  • 260326 02/10/09 16:37 adaq LHRS Q3 alarm
  • 260325 02/10/09 16:35 R. Michaels R-HRS DAQ restored
  • 260323 02/10/09 16:27 Kalyan BigBite shower x vs y
  • 260321 02/10/09 16:10 Jin Verbose Replay
  • 260320 02/10/09 16:06 O. Glamazdin BB Cherenkov Trigger histograms
  • 260318 02/10/09 16:01 Ole Hansen Day shift summary
  • 260315 02/10/09 15:33 V. Sulkosky Left HRS Gas Cerenkov
  • 260314 02/10/09 15:32 Kalyan summary of carbon runs without BB seive
  • 260313 02/10/09 15:32 Kalyan summary of carbon runs without BB seive
  • 260312 02/10/09 15:28 A. Camsonne Compton GSO to 1900 V
  • 260310 02/10/09 15:19 Jin Database: also changed db_BB.cer.dat
  • 260308 02/10/09 15:14 X. Zheng BB Cherenkov TDC
  • 260306 02/10/09 15:12 Kalyan Dark currents with 4uA on carbon
  • 260304 02/10/09 15:06 Kalyan Dark current with 7uA on carbon
  • 260302 02/10/09 14:54 Jin New database Dir for BB GC ADC crate map change
  • 260300 02/10/09 14:50 brads db_DL.dat updated for BB Cerenkov
  • 260299 02/10/09 14:45 Kalyan Dark currents with 6uA on carbon
  • 260296 02/10/09 14:29 Ole Hansen Analyzer updated
  • 260294 02/10/09 14:27 Kalyan Dark currents on BB , 5uA on carbon
  • 260291 02/10/09 14:18 brads BB Cerenkov flash ADC runs: 1244--1252
  • 260290 02/10/09 14:16 brads All BB Cerenkov PMTs back into the regular v792 ADCs
  • 260289 02/10/09 14:10 Ole Hansen Controlled access done
  • 260288 02/10/09 14:10 brads BB Sieve is OUT
  • 260287 02/10/09 13:53 flay BB Cerenkov ADCs
  • 260285 02/10/09 13:49 Ole Hansen Controlled access
  • 260284 02/10/09 13:36 Jin Preliminary Check of Coin Peak and BB Track Reconstruction of H2 Run
  • 260281 02/10/09 12:50 V. Sulkosky Pion Rejector Layer 1 Checkout
  • 260278 02/10/09 12:13 Ole Hansen Q3 unstable
  • 260269 02/10/09 12:12 Ole Hansen Carbon optics runs
  • 260266 02/10/09 11:59 kalyan offline deadtime for run 1249
  • 260265 02/10/09 11:59 kalyan offline deadtime for run 1249
  • 260260 02/10/09 11:31 brads BB Cerenkov PMTs 1--16 in flash ADC
  • 260259 02/10/09 11:29 A. Camsonne Compton photon rates ok
  • 260255 02/10/09 11:20 Ole Hansen Carbon target
  • 260254 02/10/09 11:17 Kalyan BigBite thresholds
  • 260253 02/10/09 11:16 Ole Hansen Raster settings
  • 260248 02/10/09 10:56 A. Camsonne Made backup copy of /usr/compton/bin
  • 260247 02/10/09 10:53 brads New TWOARM_FBADC coda config
  • 260246 02/10/09 10:51 Ole Hansen HRS ready
  • 260243 02/10/09 10:45 Moffit Target to BeO position
  • 260240 02/10/09 10:33 Ole Hansen Hall work finished
  • 260239 02/10/09 10:29 xqian Increase of the BigBite Wire Chamber Threshold
  • 260238 02/10/09 10:20 Jack RG-8 cable from floor to Left Stack
  • 260232 02/10/09 10:01 Yi Qiang NMR frequency sweep after field rotation: 40%
  • 260230 02/10/09 09:54 Yi Qiang Pumping direction changed to Transverse+
  • 260228 02/10/09 09:48 Jin Spectrums of He3 Running
  • 260227 02/10/09 09:44 brads BB Sieve is IN
  • 260224 02/10/09 09:37 V. Sulkosky Checked Production Runs (1202 to 1230)
  • 260221 02/10/09 09:14 A. Camsonne Compton central crystal additionnal checks
  • 260220 02/10/09 09:06 V. Sulkosky Re: S2m Pedestals Updated on a-onl Account
  • 260219 02/10/09 08:42 Yoomin Owl Shift Summary
  • 260218 02/10/09 08:41 A. Deur Lasers off
  • 260217 02/10/09 08:35 fansler/ewin spectrometer/polarity
  • 260216 02/10/09 08:32 A. Camsonne Compton scaler checked
  • 260212 02/10/09 07:31 A. Tobias Laser Absorption Spectrum Owl Shift
  • 260209 02/10/09 07:04 brads BB Cerenkov PMT 20's ADC signal cable not well plugged in
  • 260208 02/10/09 06:51 Yoomin Controlled Access
  • 260204 02/10/09 06:28 brads +1 MeV energy shift to Hall A when Hall C came up.
  • 260203 02/10/09 06:18 A. Tobias Target: NMR for Long. 3He pol. = 48.9% at 6:13AM Feb.10, 2009
  • 260199 02/10/09 06:06 Yoomin L-HRS momentum changes 1.2->1.0GeV
  • 260197 02/10/09 05:53 Yoomin CODA reboot
  • 260189 02/10/09 04:56 Yoomin BB HV trip
  • 260188 02/10/09 04:50 A. Camsonne Compton photon rate
  • 260186 02/10/09 04:46 Yoomin Beam back
  • 260185 02/10/09 04:44 A. Tobias Target: Moved back to Pol. 3He Target
  • 260184 02/10/09 04:38 A. Tobias Target: Move to Empty Position at 4:18AM
  • 260183 02/10/09 04:22 Yoomin Beam off for a while
  • 260181 02/10/09 04:14 Navaphon dead channel BB MWDC V2P?
  • 260172 02/10/09 04:08 Owl Shift Re Restart Analyzer before each BB replay
  • 260165 02/10/09 03:26 Yoomin spot++
  • 260164 02/10/09 03:19 Yoomin Charge Asymmetry stripchart
  • 260157 02/10/09 02:30 A. Tobias Target: NMR for Long. 3He pol. = 49.2% at 2:20AM Feb.10, 2009
  • 260156 02/10/09 02:29 Yoomin Beam half wave plate out
  • 260149 02/10/09 01:39 zhangyi Feb9 Swing shift summary (late entry)
  • 260148 02/10/09 01:39 zhangyi Feb9 Swing shift summary (late entry)
  • 260143 02/10/09 00:59 Jin Temp. Golden Run for He3 running
  • 260140 02/10/09 00:51 A. Tobias Target: Laser Absorption Spectrum Owl Shift
  • 260139 02/10/09 00:47 A. Tobias Target: Cell Temperatures Snapshot
  • 260138 02/10/09 00:46 A. Tobias Target: Oven Temperature Snapshot
  • Monday

  • 260130 02/09/09 23:50 flay BB Cerenkov ADCs, run# 1204
  • 260129 02/09/09 23:49 zhangyi halla tools
  • 260122 02/09/09 23:06 A. Camsonne Compton photon rate odd
  • 260121 02/09/09 22:57 flay BB Cerenkov TDCs
  • 260117 02/09/09 22:38 zhangyi raster on BeO
  • 260115 02/09/09 22:34 brads MCC reports a request for controlled access in Hall A (but it wasn't us)
  • 260113 02/09/09 22:26 Shahinyan Target move
  • 260112 02/09/09 22:24 brads Controlled access finished at 6:30pm (late entry)
  • 260110 02/09/09 22:21 brads spot++ fix for coincidence daq running (late entry)
  • 260108 02/09/09 21:59 Shahinyan Target move
  • 260107 02/09/09 21:42 flay BB Cerenkov ADCs
  • 260106 02/09/09 21:25 Jin Carbon Foils are correct in Online Replay
  • 260105 02/09/09 21:23 Jin Carbon Foils are correct in Online Replay
  • 260104 02/09/09 21:22 Yi Qiang NMR of Longitudinal pumping After rotation: 52.8%
  • 260103 02/09/09 21:12 zhangyi raster size and rates on he3
  • 260102 02/09/09 21:06 Yi Qiang Field rotated to Longitudinal +
  • 260100 02/09/09 21:02 flay BB Cerenkov
  • 260097 02/09/09 20:58 zhangyi rates on 2x2 he3 target
  • 260094 02/09/09 20:53 V. Sulkosky Run Database
  • 260093 02/09/09 20:52 Shahinyan Target move
  • 260092 02/09/09 20:52 Yi Qiang NMR of transverse pumping before rotation: 54.9%
  • 260090 02/09/09 20:46 R. Michaels re: flakey L1A
  • 260088 02/09/09 20:45 zhangyi 6x7(mcc unit), 2 uA on Ref. cell
  • 260086 02/09/09 20:38 zhangyi 4x4, 2uA rates on Ref. cell
  • 260083 02/09/09 20:31 zhangyi rates & raster size on Ref. cell
  • 260080 02/09/09 20:23 Shahinyan Target move
  • 260079 02/09/09 20:19 zhangyi raster size 2x2.6 (4x4) in MCC unit
  • 260075 02/09/09 20:13 zhangyi rates & raster size in BeO, 2uA
  • 260065 02/09/09 20:09 zhangyi raster size& rates
  • 260062 02/09/09 19:59 Shahinyan Target move
  • 260061 02/09/09 19:54 Yi Zhang last raster size entry is looking at a wrong run
  • 260058 02/09/09 19:40 zhangyi raster size 4x4
  • 260056 02/09/09 19:33 zhangyi rates on 2uA
  • 260054 02/09/09 19:30 zhangyi target -> He3
  • 260053 02/09/09 19:24 zhangyi spot++ in BeO
  • 260048 02/09/09 19:06 V. Sulkosky No Extra Synch Events for Run 1187 and 1188
  • 260047 02/09/09 19:02 V. Sulkosky Re: L1A Signal at left HRS TS
  • 260046 02/09/09 18:59 zhangyi target -> BeO
  • 260045 02/09/09 18:46 Yi Qiang First NMR after Oven Temperature set to 240
  • 260044 02/09/09 18:34 Yi Qiang Oven Temperature set to 240
  • 260040 02/09/09 17:49 O. Glamazdin Moller results
  • 260036 02/09/09 17:42 Kalyan Offline deadtime for the run 1178
  • 260035 02/09/09 17:07 Yi Qiang EPR with Transverse Pumping
  • 260034 02/09/09 17:03 brads BB MWDC gas system pressures retuned
  • 260032 02/09/09 16:03 Yi Qiang NMR Measurement at Empty with Transverse pumping, polarization is not calibrated
  • 260031 02/09/09 16:01 gilman shift summary - Feb 9 day
  • 260029 02/09/09 16:00 V. Sulkosky Re: Synch check was for run 1184
  • 260028 02/09/09 15:58 V. Sulkosky Synchronization Check of Left HRS ROCs and BigBite
  • 260027 02/09/09 15:48 xqian carbon foils
  • 260026 02/09/09 15:34 xqian This is what the spectrum should be
  • 260025 02/09/09 15:16 xqian There are some noises.
  • 260024 02/09/09 14:59 posik TDC vs ADC, buffer vs unbuffer
  • 260023 02/09/09 14:53 posik TDC vs ADC discrepancy
  • 260022 02/09/09 14:44 posik TDC vs ADC, Buffer and nonBuffer
  • 260021 02/09/09 13:54 Yi Qiang NMR Measurement at Empty with Transverse pumping, polarization is not calibrated
  • 260020 02/09/09 12:59 gilman partial shift summary Feb 9 day
  • 260019 02/09/09 12:34 V. Sulkosky Helicity from HRS
  • 260010 02/09/09 12:10 R. Michaels re: helicity
  • 260009 02/09/09 12:09 Yi Qiang NMR Measurement at Empty with Transverse pumping, polarization is not calibrated
  • 260002 02/09/09 09:40 R. Michaels empty target
  • 259999 02/09/09 09:21 R. Michaels re: helicity
  • 259995 02/09/09 09:06 R.Michaels web guides for 2-DAQ
  • 259993 02/09/09 08:53 R. Michaels re: helicity synch
  • 259991 02/09/09 08:40 mcanan BB wire hit
  • 259989 02/09/09 08:38 R. Michaels 304 Torr.
  • 259983 02/09/09 08:21 R. Michaels N2 to 207 Torr
  • 259980 02/09/09 08:06 A. Tobias Target: Reference Cell at 104 Torr N2 gas.
  • 259977 02/09/09 08:02 Yoomin Owl Shift Summary
  • 259976 02/09/09 07:58 A. Tobias Target: Laser Absorption Spectrum Owl Shift
  • 259975 02/09/09 07:56 yoomin BB PSum TDC25
  • 259973 02/09/09 07:43 A. Tobias Target: Reference Cell Evacuated ... 009 Torr / -013 psig
  • 259972 02/09/09 07:41 A. Tobias Target: Reference Cell at 2 psig = 818 Torr 3He just before evacuation
  • 259969 02/09/09 07:22 A. Tobias Target: Reference Cell at 2 psig 3He gas.
  • 259966 02/09/09 06:57 A. Tobias Target: Moved to Reference Cell with 17 psig 3He gas.
  • 259965 02/09/09 06:42 Yoomin BB trip
  • 259964 02/09/09 06:25 Jin Helicity Signal Problem
  • 259962 02/09/09 06:17 A. Tobias Target: Reference Cell at 16 psig 3He gas (target still in empty position)
  • 259961 02/09/09 06:14 A. Tobias Target: Moved to Empty Target
  • 259960 02/09/09 06:01 A. Tobias Target: Reference Cell at 50 psig.
  • 259959 02/09/09 06:00 Jin run data base hasn't been updated
  • 259953 02/09/09 05:40 A. Tobias Target: Reference Cell at 77 psig.
  • 259950 02/09/09 05:25 A. Tobias Target: Reference Cell at 106 psig
  • 259948 02/09/09 05:23 Yoomin Raster size test with ref cell(3He)
  • 259947 02/09/09 05:17 Yoomin spot++
  • 259941 02/09/09 04:45 Jin BB Chamber HV Trip
  • 259938 02/09/09 04:42 A. Camsonne Compton electron detector dak current
  • 259933 02/09/09 04:28 Yoomin Beam check with BeO
  • 259931 02/09/09 04:26 Yoomin spot++
  • 259914 02/09/09 03:55 A. Tobias Target: Moved to Empty & back to BeO
  • 259913 02/09/09 03:55 A. Camsonne HAll A Accelerator
  • 259912 02/09/09 03:03 A. Camsonne Compton rates seems OK, but odd behavior of GSO
  • 259911 02/09/09 02:29 adaq BPM snapshot
  • 259898 02/09/09 01:22 Yoomin Update from MCC
  • 259897 02/09/09 01:09 Yoomin Last HARP result
  • 259894 02/09/09 00:56 Yoomin beam spot
  • 259889 02/09/09 00:08 A. Camsonne Spot quick fix for coincidence
  • 259887 02/09/09 00:02 gilman shift summary - Feb 8 swing
  • Sunday

  • 259876 02/08/09 21:46 Yi Qiang NMR Measurement took at Optics position with Transverse pumping, polarization is not calibrated !!!!!!
  • 259875 02/08/09 21:25 Yi Qiang NMR Measurement took at Optics position with Transverse pumping, polarization is not calibrated !!!!!!
  • 259874 02/08/09 21:17 Yi Qiang Fiber configuration for Longitudinal Pumping
  • 259873 02/08/09 21:10 Yi Qiang Rotated to transverse pumping
  • 259871 02/08/09 20:46 V. Sulkosky Re: Scaler gates setup for the L-HRS in 1-DAQ Mode
  • 259870 02/08/09 20:36 V. Sulkosky Re: Left HRS Triggers for Normalization Scalers
  • 259869 02/08/09 20:32 A. Rakhman Target Moved
  • 259868 02/08/09 20:32 V. Sulkosky Re: Checked Timing on Left HRS for Twoarm_FB Mode
  • 259867 02/08/09 20:30 gilman MCC asks for BeO target
  • 259856 02/08/09 19:54 R. Michaels TWOARM_FB
  • 259851 02/08/09 19:39 Kalyan Coincidence trigger setup
  • 259849 02/08/09 19:21 V. Sulkosky Reconnected Trigger Bit and Latch Pattern Cables to TDC for Left HRS
  • 259848 02/08/09 19:17 brads BB Cerenkov ADC gate timing restored after trigger work
  • 259846 02/08/09 18:51 D. Parno Compton FADC now running at 30 Hz
  • 259845 02/08/09 18:25 A. Camsonne Compton work
  • 259844 02/08/09 18:25 A. Camsonne Compton work
  • 259841 02/08/09 17:50 brads Tech: ethane/argon gas pressure
  • 259840 02/08/09 17:40 Yi Qiang Variac is back in
  • 259839 02/08/09 17:36 brads MWDC bubblers adjusted again
  • 259838 02/08/09 17:15 brads Left arm VDC HVs tripped, can't restore
  • 259836 02/08/09 16:46 Kalyan ADC Gates and front-end busy in the scalers
  • 259835 02/08/09 16:43 Kalyan BB scalers gates problem fixed
  • 259829 02/08/09 16:28 Yi Qiang More oven work
  • 259828 02/08/09 16:21 gilman pass change
  • 259826 02/08/09 16:00 Holmstrom Day Shift Summary
  • 259823 02/08/09 15:48 Jin physics replay configured for 1 pass running
  • 259820 02/08/09 15:39 Yi Qiang Replaced the oven temperature controller
  • 259818 02/08/09 15:33 Holmstrom Spot++ pics
  • 259816 02/08/09 14:58 Holmstrom Hall A tools
  • 259813 02/08/09 14:52 W. Troth Target: Laser Absorption Spectrum Day Shift
  • 259812 02/08/09 14:51 Kalyan BB scalers gates problem
  • 259811 02/08/09 14:47 W. Troth Target: Latest NMR at 2:35PM was 44.5%
  • 259808 02/08/09 14:18 W. Troth Oven Temperature Drop
  • 259805 02/08/09 14:03 posik prescale mistype
  • 259799 02/08/09 13:09 posik prescale change
  • 259795 02/08/09 12:40 flay Summary of runs 1089-1096 MPS
  • 259784 02/08/09 11:56 flay MPS
  • 259783 02/08/09 11:55 flay MPS ROC 5, ROC 8
  • 259782 02/08/09 11:45 flay MPS ROC 5 ROC 8
  • 259779 02/08/09 11:43 W. Troth Target: Latest NMR at 11:35AM was 45.0%
  • 259778 02/08/09 11:31 posik roc5 power cycled
  • 259774 02/08/09 11:02 posik roc5 reboot
  • 259773 02/08/09 10:59 posik roc5 error
  • 259767 02/08/09 09:23 W. Troth Oven Temperature Drop
  • 259764 02/08/09 09:08 posik end run 1116
  • 259763 02/08/09 08:58 W. Troth Target: Latest NMR at 8:35AM was 46.5%
  • 259755 02/08/09 08:07 Yoomin L-HRS mometum 0.6 -> 1.6 GeV
  • 259753 02/08/09 08:00 Yoomin Owl Shift Summary
  • 259741 02/08/09 06:05 A. Tobias Target: Laser Absorption Spectrum Owl Shift
  • 259740 02/08/09 06:01 A. Tobias Target: Latest NMR at 5:35AM was 46.4%
  • 259731 02/08/09 04:25 Yoomin Strange BB ADC sum
  • 259718 02/08/09 03:56 Yoomin Raster Size Test for production
  • 259713 02/08/09 03:46 Yoomin spot++ 20114
  • 259701 02/08/09 03:13 brads Updated cer.cfg file in onlana directory
  • 259700 02/08/09 03:11 Yoomin spot++ 20111
  • 259699 02/08/09 03:07 A. Tobias Target: 3He polarization about 45.9% Screen Shot
  • 259696 02/08/09 03:03 Yoomin spot++ 20110
  • 259694 02/08/09 02:58 Yoomin spot++ 20109
  • 259689 02/08/09 02:48 adaq spot++ 20108
  • 259684 02/08/09 02:41 A. Tobias Target: Started NMR FS Control Program
  • 259681 02/08/09 02:30 A. Tobias Target: Moved to Long. Pol. 3He Position
  • 259673 02/08/09 02:01 Yi Qiang Oven temperature stablized
  • 259672 02/08/09 02:01 A. Tobias Target: Reference Cell at 105 psig 3He gas.
  • 259669 02/08/09 01:59 Yoomin No empty histogram
  • 259666 02/08/09 01:45 A. Tobias Target: Reference Cell at 75 psig 3He gas.
  • 259663 02/08/09 01:41 Kalyan Re:Empty histograms in replay
  • 259660 02/08/09 01:32 A. Tobias Target: Reference Cell at 45 psig 3He gas.
  • 259657 02/08/09 01:30 Kalyan Shower and preshower check
  • 259655 02/08/09 01:21 A. Tobias Target: Reference Cell at 15 psig 3He gas.
  • 259651 02/08/09 01:18 Yoomin Empty histograms in replay
  • 259650 02/08/09 01:17 Yoomin BB.cer.a[5]:BB.sum.rt not found
  • 259648 02/08/09 01:06 A. Tobias Target: Evacuate Reference Cell & Fill to 715 Torr (0 psig) 3He gas.
  • 259646 02/08/09 00:58 Yoomin N2 ref cell pressure curve done
  • 259643 02/08/09 00:49 Yoomin Prescal changed for BB
  • 259640 02/08/09 00:45 A. Tobias Target: Reference Cell at 602Torr.
  • 259637 02/08/09 00:35 A. Tobias Target: Reference Cell at 403Torr
  • 259636 02/08/09 00:33 A. Tobias Target: Oven Temperature dropped to about 210C
  • 259631 02/08/09 00:26 flay TDC cerenkov with t2 cut
  • 259630 02/08/09 00:21 xqian Hit effciency
  • 259623 02/08/09 00:06 brads LHRS CODA died
  • 259622 02/08/09 00:05 LeRose Shift Summary (swing)
  • Saturday

  • 259619 02/07/09 23:56 Moffit ROC5 Errors in buffered mode
  • 259604 02/07/09 23:15 posik prescales changed
  • 259597 02/07/09 22:57 Yunxiu Ye Laser_abs_spectr_020709_swing
  • 259590 02/07/09 22:37 posik 3x4 mcc rates
  • 259589 02/07/09 22:35 Yunxiu Ye Reference target is in
  • 259588 02/07/09 22:34 LeRose Spot++
  • 259586 02/07/09 22:29 brads BB Cerenkov work: TDC delay cables fixed, Disc. output widths increased
  • 259582 02/07/09 22:13 V. Sulkosky Left HRS PID Plots
  • 259581 02/07/09 22:06 LeRose Power Permit
  • 259580 02/07/09 22:01 Yunxiu Ye Target optics
  • 259577 02/07/09 21:52 Moffit Ouch... camac fuse popped again. Replaced with the correct fuse.
  • 259576 02/07/09 21:40 V. Sulkosky Pedestal Suppression File Updated for S2m
  • 259575 02/07/09 21:35 xqian Chamber background
  • 259574 02/07/09 21:31 Yunxiu N2 pressure 006 tott
  • 259573 02/07/09 21:31 posik mwdc oddity
  • 259572 02/07/09 21:25 V. Sulkosky Re: S2m Pedestals updated in Podd Database
  • 259571 02/07/09 21:20 V. Sulkosky Modified Left HRS Replay Cut File
  • 259570 02/07/09 21:20 Yi Qiang NMR Frequency sweep procedure is posted on wiki
  • 259569 02/07/09 21:01 Yi Qiang EPR calibration with only two COMET lasers
  • 259568 02/07/09 20:47 xqian acceptance is ok
  • 259567 02/07/09 20:36 Yi Qiang EPR Calibration for smantha
  • 259558 02/07/09 19:46 Moffit Replaced 5A fuse on BBcamac crate.
  • 259557 02/07/09 19:13 V. Sulkosky Target Distributions (?)
  • 259556 02/07/09 19:04 V. Sulkosky Theta_tg Distribution
  • 259555 02/07/09 18:49 posik runs 1084-1086 no good
  • 259554 02/07/09 18:43 LeRose Controlled Access
  • 259553 02/07/09 18:43 Moffit BB CAMAC crate down again.
  • 259551 02/07/09 18:39 LeRose DAQ Problems
  • 259542 02/07/09 17:47 LeRose BCM Readout????
  • 259537 02/07/09 17:38 posik bad runs 20082,20083,1080,1081
  • 259528 02/07/09 17:17 Yunxiu Ye changing N2 presure to 605
  • 259515 02/07/09 16:38 Moffit S14 MWDC
  • 259512 02/07/09 16:36 Moffit MWDC HV for 800 torr N2, 14uA
  • 259511 02/07/09 16:33 Yi Qiang Target Move to Reference Cell: N2 816 torr
  • 259510 02/07/09 16:26 LeRose Spott++
  • 259503 02/07/09 16:06 Holmstrom Day Shift Summary
  • 259501 02/07/09 15:42 Moffit Moved T4 from MLU output 1 to output 2
  • 259500 02/07/09 15:41 Yi Qiang Access for BeO camera
  • 259499 02/07/09 14:49 Troth Hall A Tools
  • 259498 02/07/09 14:46 Holmstrom BeO camara Problem
  • 259497 02/07/09 14:43 Holmstrom BB wire Chambers HV-On
  • 259496 02/07/09 14:39 E.Chudakov Moller results
  • 259495 02/07/09 13:25 V. Sulkosky Updated Scripts for db_run.dat database
  • 259494 02/07/09 13:07 flay Cerenkov plots
  • 259493 02/07/09 12:29 posik updated cer.cfg
  • 259484 02/07/09 11:12 R. Michaels modified L-HRS datamon
  • 259483 02/07/09 11:04 R. Michaels T4 problem
  • 259482 02/07/09 10:30 Holmstrom BB HV - OFF
  • 259481 02/07/09 10:26 V. Sulkosky Spot++ Fixed
  • 259480 02/07/09 10:25 flay BPM Harp Scan Calibration
  • 259479 02/07/09 10:25 flay BPM Harp Scan Calibration
  • 259478 02/07/09 10:10 Holmstrom Target moved to Empty
  • 259469 02/07/09 09:34 flay BB Pedistals
  • 259468 02/07/09 09:27 Holmstrom Temps coming back
  • 259467 02/07/09 09:15 Holmstrom Oven Temp dip
  • 259466 02/07/09 09:12 flay BB pedistals
  • 259451 02/07/09 08:10 Yoomin Owl Shift Summary
  • 259447 02/07/09 07:48 adaq BPM/BCM
  • 259446 02/07/09 07:29 Yoomin MCC have difficulty in moving BPM_Ys
  • 259445 02/07/09 07:24 Yoomin Current were 2uA.
  • 259444 02/07/09 07:15 Yoomin spot check
  • 259443 02/07/09 07:00 Yoomin BB HV trip
  • 259442 02/07/09 06:49 Yoomin beam_xy 20069
  • 259441 02/07/09 06:45 yawei move target to optics with hole
  • 259438 02/07/09 06:32 Yoomin beam_xy 20068
  • 259437 02/07/09 06:31 adaq beam_xy 20067
  • 259436 02/07/09 06:29 yawei rates for raster 6x6 MCC
  • 259431 02/07/09 06:21 Yoomin spot++ doesn't work
  • 259430 02/07/09 06:21 yawei rates for raster 5x5 MCC
  • 259427 02/07/09 06:13 yawei rates for raster 3x4 MCC
  • 259421 02/07/09 05:58 yawei target moved to reference cell
  • 259420 02/07/09 05:55 jpchen Tech: Target camera light out.
  • 259412 02/07/09 05:23 Yoomin Beam permit
  • 259403 02/07/09 02:56 brads Target camera view is very unclear
  • 259402 02/07/09 02:22 Yoomin Controlled access asked.
  • 259401 02/07/09 02:19 A. Camsonne Photon compton rates good
  • 259400 02/07/09 02:18 Yoomin BPM check
  • 259399 02/07/09 02:15 yawei target moved to optics
  • 259398 02/07/09 02:12 brads Set thresholds on BigBite
  • 259397 02/07/09 02:04 Yoomin Beam Permit
  • 259394 02/07/09 02:02 brads 'bigbite_buf' SQL database slightly corrupted, fixed
  • 259391 02/07/09 01:52 V. Sulkosky Re: Scalers in End-of-run file for d2n
  • 259389 02/07/09 01:38 Target Group Adjustment of Q-wave plates for Longitudinal Lasers
  • 259382 02/07/09 01:13 yawei target moved to empty
  • 259380 02/07/09 01:06 brads BB Cerenkov PMTs 2 and 20 look very poor
  • 259379 02/07/09 01:00 brads Pedestals for the linear gated BB Cerenkov signals
  • 259377 02/07/09 00:51 brads BB Shower threshold calibration constant
  • 259376 02/07/09 00:49 brads re: BB Cerenkov cabling and ADC map clarification
  • 259375 02/07/09 00:48 Moffit Added BB Cerenkov to replay and DB
  • 259371 02/07/09 00:25 brads/bogdan BB Cerenkov cabling and ADC map changes
  • 259370 02/07/09 00:16 bogdanw shift summary
  • 259368 02/07/09 00:15 brads He3 bottle has been opened (no leaks observed)
  • Friday

  • 259360 02/06/09 23:15 Kalyan Scaler gates setup for the L-HRS
  • 259353 02/06/09 22:17 brads Reference cell gas panel on upper monitor in middle room
  • 259352 02/06/09 21:56 R. Michaels re:deadtime vs rate
  • 259351 02/06/09 21:20 A. Camsonne RICH fast clear discriminator removed
  • 259350 02/06/09 20:35 Kalyan Dead time study for buffered and unbuffered mode on the BigBite DAQ
  • 259349 02/06/09 20:23 brads re: Lots of bubbling through the INflow MWDC bubbler
  • 259348 02/06/09 20:21 Yi/Jin Hall A is Locked for Laser work
  • 259347 02/06/09 20:09 brads ~320 ns of delay added to Cerenkov TDC signals
  • 259338 02/06/09 20:00 Jin Target data backup for d2n
  • 259321 02/06/09 19:32 V. Sulkosky Re: Start/End Run Script Status for d2n
  • 259319 02/06/09 19:27 V. Sulkosky Re: E06-014 (d2n) MySQL Database
  • 259314 02/06/09 19:12 brads BB Sieve is OUT
  • 259279 02/06/09 17:14 Jin Spin signal set to enable all scalars working
  • 259278 02/06/09 16:38 V. Sulkosky Left HRS Triggers for Scalers and TDC
  • 259277 02/06/09 16:37 Kalyan Removed delay for TDC gate on ROC8
  • 259274 02/06/09 16:19 V. Sulkosky S2m Analyzer database updated
  • 259273 02/06/09 16:19 V. Sulkosky S2m Analyzer database updated
  • 259271 02/06/09 16:12 jpchen/yiqia pressure gauge/cooling jet flow
  • 259270 02/06/09 16:10 V. Sulkosky No EDTM Signal on Left HRS
  • 259269 02/06/09 16:06 V. Sulkosky Moved S2m Left Side ADC Channels
  • 259267 02/06/09 15:53 V. Sulkosky GC HV Channel Moved
  • 259266 02/06/09 15:45 folts checklist
  • 259206 02/06/09 11:52 R. Michaels R-HRS DAQ for parasitic pleasures.
  • 259200 02/06/09 11:29 Moffit Reconnected BB ground
  • 259192 02/06/09 10:39 Moffit oddity in transversity's GUI configuration for detectorBB.cfg
  • 259191 02/06/09 10:32 R. Michaels DAQ status, things done and to do
  • 259187 02/06/09 10:22 Kalyan New analyzer database created for d2n expt
  • 259150 02/06/09 08:26 atrig Trigger-Download
  • 259148 02/06/09 05:16 Target Group Laser alignment done, CAN system disabled
  • 259139 02/06/09 03:04 lrichard IHOP
  • 259138 02/06/09 02:38 ALL present IHOP
  • 259137 02/06/09 01:51 lrichard IHOP
  • 259136 02/06/09 01:26 All present IHOP
  • 259135 02/06/09 01:15 Jin Online Analysis Software is Setup for d2n
  • 259134 02/06/09 00:11 V. Sulkosky E06-014 (d2n) MySQL Database
  • 259133 02/06/09 00:07 V. Sulkosky Start/End Run Script Status for d2n
  • Thursday

  • 259129 02/05/09 23:47 Jin run database modification
  • 259124 02/05/09 22:24 Yi/chiranjib Laser alignment status
  • 259122 02/05/09 22:08 Jin Target: Stepper motors driving laser 1/4Wave plates are OFF
  • 259120 02/05/09 21:09 Moffit DAQ: Bigbite CAMAC crate revived
  • 259118 02/05/09 20:16 Jin Hall A going Laser control
  • 259117 02/05/09 20:15 Jin Hall A going Laser control
  • 259097 02/05/09 18:05 Katich d2n target installation update
  • 259096 02/05/09 17:43 R. Michaels gate for TDC slot 9 roc 4
  • 259094 02/05/09 17:41 R. Michaels ROC4 issues
  • 259085 02/05/09 16:57 zhangyi target laser fibers switched from vertical to longitudinal.
  • 259083 02/05/09 16:31 V. Sulkosky RICH HV Off
  • 259081 02/05/09 16:27 V. Sulkosky A1 HV's Turned Off
  • 259073 02/05/09 14:39 Joe/Chiranji he-3 cell in / ladder in
  • 259069 02/05/09 13:54 Chiranjib COMPASS survey of the field DONE
  • 259060 02/05/09 11:41 Yi Qiang Current settings for compass measurement
  • 259055 02/05/09 10:10 Moffit DAQ: Bigbite CAMAC power supply removed
  • 259054 02/05/09 10:04 adev BigBite WireChamber Threshold change!
  • 259052 02/05/09 09:32 Yi Qiang Lasers are LOCKED
  • 259051 02/05/09 09:20 NMR Cross Calibration for Vertical pumping
  • 259050 02/05/09 09:05 Yi Qiang AFP Loss test
  • 259049 02/05/09 09:04 Yi Qiang AFP Loss test
  • 259048 02/05/09 08:42 Yi Qiang Internal Temperature Test with Vertical Pumping
  • 259046 02/05/09 07:58 riordan Owl Shift Summary
  • 259045 02/05/09 07:48 Jin/YI EPR calibration, both pumping direction
  • 259044 02/05/09 07:20 ycwang RICH calibration runs
  • 259043 02/05/09 07:19 Jin Cosmic run for Bigbtie Shower Calibration
  • 259037 02/05/09 07:02 riordan Stop beam, going to controlled access
  • 259035 02/05/09 06:31 Jin Huang Summary : Bigbite Optics Sieved Runs @ 5pass beam
  • 259032 02/05/09 06:22 Jin Further Backup/Remove Online Root Files
  • 259030 02/05/09 06:21 kelleher Laser Spectrum for Owl Shift
  • 259026 02/05/09 05:16 riordan LHRS to 0.6 GeV/c
  • 259023 02/05/09 04:20 riordan spot++, 6x6 raster
  • 259020 02/05/09 04:12 kelleher Target Moved to 3He position
  • 259019 02/05/09 04:06 riordan Target Change, Raster on
  • 259004 02/05/09 02:57 riordan LHRS to 1 GeV/c and Raster Off
  • 259002 02/05/09 02:16 riordan Hall A Tools
  • 259001 02/05/09 02:14 riordan spot++
  • 258994 02/05/09 01:43 riordan Changing LHRS momentum setting
  • 258989 02/05/09 01:25 brads Zeroed DC offsets on BB Cerenkov linear gate unit
  • 258986 02/05/09 01:16 A. Camsonne RICH pedestal run 5875
  • 258985 02/05/09 01:15 brads Plugged the latch pattern ribbon cable back (effect: no evtypebits var for runs 5868--5872), now fixed
  • 258983 02/05/09 01:13 A. Camsonne RICH ADC limit set to 480
  • 258981 02/05/09 01:11 brads BB Sieve is IN
  • 258976 02/05/09 00:49 riordan Access for BigBite Sieve
  • 258974 02/05/09 00:34 bogdanw shift summary
  • 258972 02/05/09 00:14 brads Cable on 4518 was knocked loose and restored during Access
  • 258971 02/05/09 00:07 brads Run 5869 is spot++ run w/ BeO target, not production
  • 258970 02/05/09 00:06 brads Run 5868 is pedestal run, not production
  • Wednesday

  • 258966 02/04/09 23:42 brads BB Cerenkov PMT 5 HV restored to 1215V
  • 258965 02/04/09 23:35 brads Lots of bubbling through the INflow MWDC bubbler
  • 258964 02/04/09 23:34 bogdanw spot++ for run #5869
  • 258961 02/04/09 23:30 brads/bogdan 8 BB Cerenkov PMTs now through linear gate
  • 258959 02/04/09 23:24 zhihong Ye Target moved from he3 to optics
  • 258952 02/04/09 20:41 Zhihong Ye Target Laser absorption spectrum and Vertical polarization
  • 258935 02/04/09 17:31 A. Camsonne Compton HV set to 1550 V
  • 258934 02/04/09 17:24 adaq spot++ run 5861
  • 258930 02/04/09 16:00 X. Zheng day shift summary
  • 258927 02/04/09 14:49 X. Zheng Door handle
  • 258924 02/04/09 13:41 Katich clarification
  • 258923 02/04/09 12:53 narayan Laser Absorption Spectrum
  • 258911 02/04/09 11:37 Katich d2n 3He Cell Mounted
  • 258910 02/04/09 11:22 X. Zheng spot++ (run 5857)
  • 258907 02/04/09 10:11 yawei left arm VDC efficiency, run#5854
  • 258904 02/04/09 10:00 V. Sulkosky Gas Cerenkov ADC #6 Status
  • 258900 02/04/09 07:56 Carlos Owl shift summary
  • 258896 02/04/09 07:28 kelleher Owl Shift Laser Spectrum
  • 258887 02/04/09 06:57 Carlos spot++
  • 258886 02/04/09 06:23 ycwang nice Charge Asymmetry goes on
  • 258881 02/04/09 03:50 Jin Re: Helicity Diff
  • 258878 02/04/09 03:29 ycwang Helicity diff does not match the gold run?
  • 258877 02/04/09 03:23 Sarah Weird Histogram for Helicity Diff?
  • 258866 02/04/09 01:15 brads Controlled Access: BB Cerenkov PMT 5 sent through linear gate
  • 258865 02/04/09 01:13 Jin Beam energy lock back to normal
  • 258860 02/04/09 00:46 brads Powercycled TS12 again to fix scaler problem...
  • 258853 02/04/09 00:30 brads TS12 rebooted to fix scaler problem
  • 258847 02/04/09 00:22 brads Rebooted a bunch of ROCs to get CODA working
  • 258846 02/04/09 00:21 Kalyan CAMAC crate reboot : Thresholds set back to normal
  • 258843 02/04/09 00:19 adev BigBite WireChamber Threshold change!
  • 258837 02/04/09 00:08 bogdanw shift summary
  • Tuesday

  • 258835 02/03/09 23:38 adev BigBite WireChamber Threshold change!
  • 258833 02/03/09 23:18 Jin Beam Energy Fluctuating
  • 258832 02/03/09 23:01 Jin 100% dead time, rebooting CODA
  • 258830 02/03/09 21:40 Jin Laser Spectrum
  • 258819 02/03/09 21:03 Typo Not "target chamber RTD", should be "Pumping Chamber RTD "
  • 258818 02/03/09 21:01 adev BigBite WireChamber Measured Voltage change!
  • 258816 02/03/09 20:56 Yi Qiang Last RTD on target chamber gone...
  • 258814 02/03/09 20:37 A. Camsonne Final tune for Compton, rates are good
  • 258813 02/03/09 20:34 R. Michaels re:Lumi data
  • 258812 02/03/09 20:28 A. Camsonne Compton rates after Moeller
  • 258811 02/03/09 20:23 Katich Lumi Data Insertion Error
  • 258810 02/03/09 19:45 E.Chudakov Moller results
  • 258808 02/03/09 19:23 A. Camsonne Helicity setting
  • 258806 02/03/09 17:40 A. Camsonne Compton rates very high turning off photon detector during moeller
  • 258805 02/03/09 16:01 X. Zheng Day shift summary
  • 258803 02/03/09 15:48 adev BigBite WireChamber Measured Voltage change!
  • 258802 02/03/09 15:46 Kalyan BigBite wirechamber HV turned off
  • 258801 02/03/09 15:37 adev BigBite WireChamber Measured Voltage change!
  • 258800 02/03/09 15:37 X. Zheng beam tuning
  • 258799 02/03/09 15:30 X. Zheng beam back
  • 258798 02/03/09 14:52 X. Zheng Hall going to power permit/BeO in
  • 258797 02/03/09 14:50 averett target to BeO position
  • 258796 02/03/09 14:50 Yi Qiang We are running with 2 Comet + 1 FAP
  • 258795 02/03/09 14:49 averett cooling jets?
  • 258794 02/03/09 14:42 X. Zheng Beam will be back at 3pm
  • 258793 02/03/09 14:42 averett target temps
  • 258792 02/03/09 14:39 averett target temperatures and laser spectrum
  • 258788 02/03/09 13:17 V. Sulkosky Moved Gas Cerenkov ADC #7 Delay Cable
  • 258787 02/03/09 13:16 Yi Qiang Target polarization will be high enough after 6:00pm
  • 258786 02/03/09 13:12 Moffit Added and removed modules for ROC5 VME crate
  • 258785 02/03/09 13:08 Moffit ROC5 Changes to c792 ADCs
  • 258783 02/03/09 13:05 Yi Qiang EPR for Vertical Pumping
  • 258782 02/03/09 13:05 Yi Qiang EPR for Vertical Pumping
  • 258781 02/03/09 12:55 Yi Qiang COMET laser power test
  • 258780 02/03/09 12:53 Kalyan Helicity and MPS signal in BigBite ADC
  • 258779 02/03/09 12:39 A. Camsonne Compton electron detector on hareboot4 port 4
  • 258776 02/03/09 12:31 Yi/Chiranjib Field rotated to Vertical +
  • 258768 02/03/09 11:54 Chiranjib NMR frequency / field sweep calibration
  • 258765 02/03/09 11:46 Chiranjib/ Y Internal Temperature Test
  • 258760 02/03/09 10:50 xqian BB Wire Chamber Gas
  • 258759 02/03/09 10:23 Yi Qiang Ladder control is back
  • 258757 02/03/09 10:03 V. Sulkosky Run 5818 is a Pedestal run
  • 258756 02/03/09 10:01 Yi Qiang EPR for transverse pumping
  • 258755 02/03/09 09:54 Yi Qiang Target ladder stuck again
  • 258751 02/03/09 09:12 brads re: auto-snap
  • 258750 02/03/09 09:04 averett target spin flip suspended
  • 258745 02/03/09 07:50 Carlos Owl shift summary
  • 258744 02/03/09 07:19 Carlos spot++
  • 258743 02/03/09 07:11 kelleher auto-snap?
  • 258741 02/03/09 07:02 kelleher Owl laser absorption spectrum
  • 258734 02/03/09 03:23 ycwang synchronizaton between ROC 10 and ROC4&ROC3
  • 258731 02/03/09 02:04 ycwang Compton continues
  • 258730 02/03/09 02:01 Jin Description of Compton Symptoms
  • 258727 02/03/09 01:25 zhangyi compton doesn't work
  • 258726 02/03/09 00:15 Sarah All DAQs rebooted
  • 258723 02/03/09 00:02 bogdanw shift summary
  • Monday

  • 258718 02/02/09 23:25 Yunxiu Ye Abs_spectrum_020209_swing
  • 258706 02/02/09 16:47 brads re: BB gas bubbler history
  • 258702 02/02/09 16:01 Ole Hansen Day Shift Summary
  • 258701 02/02/09 15:40 Joe Coffee problem solved
  • 258696 02/02/09 13:55 Shift Worker Coffee Emergency
  • 258695 02/02/09 13:24 Kalyan BB gas flow in the shed looks okay
  • 258692 02/02/09 12:49 Kalyan Re: BB gas not bubbling?
  • 258691 02/02/09 12:44 Ole Hansen BB gas bubbler history
  • 258690 02/02/09 12:34 Ole Hansen spot++ second screen
  • 258689 02/02/09 12:28 Ole Hansen BB gas not bubbling?
  • 258688 02/02/09 12:28 yawei left arm TDC efficiency,run#5798
  • 258687 02/02/09 12:18 Ole Hansen spot++ screenshot
  • 258684 02/02/09 10:51 V. Sulkosky Gas Cerenkov PMT #6
  • 258683 02/02/09 10:37 Ole Hansen Beam stable now
  • 258682 02/02/09 10:32 adev BigBite WireChamber Measured Voltage change!
  • 258681 02/02/09 10:28 Ole Hansen Trippy beam
  • 258679 02/02/09 10:27 Ole Hansen BB HV back up
  • 258677 02/02/09 10:22 Ole Hansen BB wire chamber trip
  • 258676 02/02/09 10:21 adev BigBite WireChamber Measured Voltage change!
  • 258670 02/02/09 07:58 Carlos Owl shift summary
  • 258669 02/02/09 07:53 Sarah Run bad because of Happex DAQ crash
  • 258668 02/02/09 07:50 Sarah Happex CODA crash
  • 258667 02/02/09 07:48 ycwang L-arm VDC efficiency
  • 258662 02/02/09 06:52 Carlos spot++
  • 258661 02/02/09 06:46 B. Norum Compton Strip Chart
  • 258660 02/02/09 06:38 B Norum Compton Strip Chart
  • 258643 02/02/09 01:06 Yi^2 Chamber test
  • 258640 02/02/09 00:59 zhangyi BB chamber HV alarm
  • 258637 02/02/09 00:21 Yi Qiang Bigbite Chamber HV tripped during run 5782 through 5785
  • 258633 02/02/09 00:01 bogdanw shift summary
  • Sunday

  • 258630 02/01/09 23:44 Yi Qiang Rates with 12uA beam on 3He target
  • 258627 02/01/09 23:32 Yi Qiang Spot++ of run 5784 (8uA)
  • 258626 02/01/09 23:23 Yi Qiang Rates with 8uA beam on 3He target
  • 258621 02/01/09 22:57 Yi Qiang Spot++ of run 5782 (4uA)
  • 258620 02/01/09 22:56 Yi Qiang Rates with 4uA beam on 3He target
  • 258617 02/01/09 22:27 Yi Qiang Rates with 2uA beam on 3He target
  • 258612 02/01/09 21:56 Yi Qiang Rates with 1uA beam on 3He target
  • 258609 02/01/09 21:45 Yi Qiang Beam Charge Asymmetry Stripchart abnormal
  • 258603 02/01/09 21:14 A. Rakhman Beam Quality from Compton rates
  • 258592 02/01/09 19:43 A. Rakhman Compton Strip Chart
  • 258591 02/01/09 19:41 A. Rakhman BPM & BCM
  • 258590 02/01/09 19:38 A. Rakhman Traget Screen Strip Charts
  • 258589 02/01/09 19:30 A.Rakhman Laser Absorption spectrum
  • 258587 02/01/09 19:10 bogdanw spot++
  • 258581 02/01/09 17:07 bogdanw PMT #6
  • 258578 02/01/09 15:58 Ge Jin day shift summary:
  • 258575 02/01/09 13:52 ge jin hall A tools screen
  • 258574 02/01/09 13:49 ge jin spot++
  • 258571 02/01/09 13:41 yawei left arm VDC, run#5769
  • 258570 02/01/09 12:54 ycwang BB HV trip in RUN 5768
  • 258569 02/01/09 12:44 Bo Laser Absorption spectrum
  • 258553 02/01/09 10:45 ycwang RUN # 5765 stopped and 5766 is checked out
  • 258549 02/01/09 10:18 Ge Jin BB shower HV tripped
  • 258543 02/01/09 08:14 ycwang When beam tripped, the deadtime shows negative on datamon display
  • 258541 02/01/09 07:49 Carlos Owl shift summary
  • 258538 02/01/09 06:16 R. Michaels re: scaler plots
  • 258535 02/01/09 06:03 R. Michaels scaler plot problem ?
  • 258533 02/01/09 05:20 R. Michaels Absorption spectrum
  • 258532 02/01/09 04:58 R. Michaels spot++
  • 258521 02/01/09 03:45 Sarah ROC12 reboot
  • 258514 02/01/09 00:11 bogdanw shift summary